### v1

# CLAP: Unsupervised 3D Representation Learning for Fusion 3D Perception via Curvature Sampling and Prototype Learning

Abstract
Unsupervised 3D representation learning via masked-and-reconstruction with differentiable rendering is promising to reduce the labeling burden for fusion 3D perception. However, previous literature conduct pre-training for different modalities separately because of the hight GPU memory consumption. Consequently, the interaction between the two modalities (images and point clouds) is neglected during pre-training. In this paper, we explore joint unsupervised pre-training for fusion 3D perception via differentiable rendering and propose CLAP, short for Curvature sampLing and swApping Prototype assignment prediction. The contributions are three-fold. 1) To overcome the GPU memory consumption problem, we propose Curvature Sampling to sample the more informative points/pixels for pre-training. 2) We propose to use learnable prototypes to represent parts of the scenes in a common feature space and bring the idea of swapping prototype assignment prediction to learn the interaction between the two modalities. 3) To further optimize learnable prototypes, we propose an Expectation-Maximization training scheme to maximize the similarity between embeddings and prototypes, followed by a Gram Matrix Regularization Loss to avoid collapse. Experiment results on NuScenes show that CLAP achieves $300\%$ more performance gain as compared to previous SOTA 3D pre-training method via differentiable rendering. Codes and models will be released.
1 Introduction
3D perception provides understanding of the ego vehicle’s surrounding 3D space and plays an important role in downstream prediction and control tasks in autonomous driving. Single-modality 3D perception has long been studied with camera [11, 6, 40, 45, 55, 19] or LiDAR (Light-Detection-And-Ranging) [36, 50, 39, 54, 12, 1] inputs, which is less robust to changing lighting and weather conditions [30]. To compensate shortages of different modalities, research prevails on fusion 3D perception with camera images and LiDAR point clouds [27, 25, 21, 24, 20, 31, 7].
Figure 1: Unlike previous unsupervised 3D representation learning method UniPAD [53] that separately pre-train LiDAR and camera encoders (a), our proposed method CLAP conducts joint pre-training for fusion perception.
Meanwhile, labeling in 3D space is notoriously time-and-energy-consuming. Unsupervised 3D representation learning [23, 18, 5, 52, 48, 17, 57], which pre-trains backbones without any label and applies the pre-trained weight to initialize downstream models for performance improvement, shows potential to alleviate labeling burden in 3D perception. Hence, the community starts to explore pre-training for fusion 3D perception in an unsupervised manner [41, 56, 22, 53], among which UniPAD [53] shows superior performance. UniPAD [53] is a masked-reconstruction-based method and utilizes differentiable rendering as decoder during pre-training. However, during pre-training stage, the camera and LiDAR encoders are separately pre-trained without considering the interaction between the two modalities in [53]. The reason stems from the high GPU memory consumption. If all of the LiDAR points and camera pixels are considered for pre-training, the most advanced GPU up-to-date (H100) is only able to hold batchsize=1. But batchsize is an important hyperparameter for image encoder and a small one normally brings degradation to the performance, making it unfeasible for 3D pre-training.
In this paper, we explore joint unsupervised pre-training for fusion 3D perception via differentiable rendering and propose CLAP, which is shorthand for Curvature sampLing and swApping Prototype assignment prediction. To begin with, we need to sample the LiDAR points and camera pixels for pre-training in order to make joint pre-training feasible. The most direct way is to use uniform sampling but as the sample number is small compared to the raw inputs ($\sim\frac{1}{100}$), uniform sampling does not brings improvement against separate pre-training. In order to sample more informative parts of the environment for pre-training, CLAP first estimates curvature of each LiDAR point in the 3D space by taking second order derivative of the SDF (signed distance field) function. Then we compute weights proportional to the estimated curvature and sample more informative points according to these weights. As for pixel sampling, we project the LiDAR points onto image planes and assign the weights to pixel. As LiDAR point clouds are sparse, we further apply gaussian blur to broadcast the weights to nearby pixels and sample pixels according to these weights. With Curvature Sampling, we further consider how to better utilize the interaction between camera images and LiDAR point clouds to gain some level of understanding about “objectness” or “parts of objects” in an unsupervised way. Inspired by SwAV [3], we propose to use learnable prototypes to learn to represent parts of the scenes. Similar to [3], CLAP applies a swapping prototype assignment prediction loss between camera and LiDAR embeddings to learn the common feature space. To further optimize the learnable prototypes, an Expectation-Maximization training scheme is proposed. In the Expectation step, we compute the probability that each prototype is assigned to the 3D embeddings in the 3D space. Then in the Maximization step, we maximize the similarity between embeddings and prototypes by minimizing the entropy of the assignment scores. Finally, to avoid collapse of the prototype [3], a gram matrix regularization Loss is proposed to minimize the correspondence among prototypes.
Through extensive experiments on the popular autonomous driving dataset NuScenes [2], we demonstrate that CLAP is able to achieve up to $300\%$ more improvement than previous SOTA method [53].
2 Related Work
Fusion 3D Object Detection. Light-Detection-And-Ranging (LiDAR) and camera are important sensors for autonomous driving perception. Previous works mainly focus on single-modality 3D perception. For LiDAR-based 3D object detection, there are three main streams with different embedding schemes for point clouds inputs. 1) Embraced by [36, 38], point-based methods utilize point-level embeddings for 3D object detection. 2) Voxel-based methods [50, 10, 54, 1, 51, 12] voxelize the 3D scene and use sparse convolution or transformer for embedding. 3) Point-voxel-combined methods [37, 39] utilize both embeddings from 1) and 2). For camera-based 3D perception, [11, 6, 40, 45, 55, 19] embed 2D features on image plane and project these 2D features into 3D space with estimated depth. As single-modality perception can be degraded by various weather and lighting conditions [30], fusion 3D perception [7, 27, 25, 21, 24, 20, 31] starts to flourish. These methods mainly focus on the way to integrate embeddings of different modalities and use supervised pre-trained 2D and 3D backbones to achieve the SOTA performance. As labeling in 3D space is costly, we explore unsupervised 3D representation learning for fusion 3D perception.
3D Pre-training. Annotating for 3D data is notoriously time-and-energy-consuming and the emergence of unsupervised representation learning for 2D image [14, 43, 3, 13, 46, 15] provides a promising way to alleviate the annotation burden. Researcher starts to introduce unsupervised 3D representation learning into scene-level 3D point clouds [47, 26, 16, 5, 18, 23, 34, 52, 48, 17, 57], which can be divided into two contrastive-based and masked-and-reconstruction-based branches. Embraced by [47, 26, 16, 5, 18, 23, 34], contrastive-based works propose various ways to build suitable views and conduct contrastive learning to improve the performance in downstream perception task. Inspired by [15] in image domain, [52, 48, 17, 57] propose to first mask the input point clouds and pre-train the 3D encoders with a shallow decoder for reconstructing the unmasked inputs. [41, 56, 22] are pioneering works to introduce contrastive learning into fusion perception. They consider camera and LiDAR embeddings as different views of the scene and apply the contrastive loss between the two modalities. Inspired by the success of neural field in representing 3D scenes [32, 44] and previous attempts to introduce neural rendering to 3D pre-training for point clouds [49, 17, 57], UniPAD [53] proposes to use a differentiable-rendering decoder for masked-and-reconstruction pre-training and achieves SOTA performance for unsupervised 3D representation learning on fusion 3D perception. However, due to the high GPU memory consumption, UniPAD [53] is only able to separately pre-train the image and point cloud encoders and fails to utilize the interaction between modalities during pre-training. In this paper, we explore unsupervised joint pre-training for 2D and 3D backbones via differentiable rendering.
3 Method
Figure 2: The pipeline of CLAP. In order to jointly pre-train the LiDAR, camera and fusion encoders, we first embed the paired LiDAR point clouds and camera images with $f^{\text{enc}}_{\text{P}}$, $f^{\text{enc}}_{\text{I}}$ and $f^{\text{enc}}_{\text{fusion}}$. Then based on the fusion features, CLAP applies differentiable rendering to predict both depth and rgb with the SDF and RGB values of the sampled points along LiDAR/camera rays from $f^{\text{SDF}}$ and $f^{\text{RGB}}$, with which we compute loss against the observed LiDAR point cloud and camera images. To make joint pre-training feasible, we propose Curvature Sampling to sample informative parts of the 3D scene, as descriped in (a). Furthermore, we propose to use learnable prototypes to represent parts of objects and use a swapping prototype assignment prediction loss to incorporate the interaction of different modalities into pre-training. Finally, we optimize the learnable prototypes with an Expectation-Maximization scheme together with a gram matrix regularization loss.
In this section, we introduce CLAP for unsupervised 3D pre-training via differentiable rendering on fusion 3D perception. As described in Fig. 2, CLAP pre-trains the image, LiDAR and fusion encoders jointly with Neural Field Rendering. In order to make joint pre-training feasible, Curvature Sampling is proposed as shown in Fig. 2 (a). To further make use of both modalities, we propose a swapping prototype assignment prediction scheme, which is visualized in Fig. 2 (b). We first discuss the formulation and overall pipeline in Section 3.1. Then we introduce the details about neural field and differentiable rendering in Section 3.2. Finally, we describe the curvature sampling and swapping prototype assignment prediction separately in Section 3.3 and 3.4.
3.1 Formulation and Pipeline
Notations. To begin with, we denote the input image set from $N_{\text{cam}}$ cameras as =$\mathcal{I}=\{\mathbf{I}_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N_{\text{cam}}}$ and LiDAR point cloud as $\mathbf{P}\in\mathbb{R}^{N_{p}\times(3+d)}$. $H$ and $W$ are the height and width of the images and each pixel on the images has $3$ values for RGB. $N_{p}$ is the number of points in the LiDAR point cloud and each of them contains $xyz$-location and $d$ feature channels. For example, in NuScenes [2] dataset, $d=2$ represents the intensity and timestamp of each point and there are $N_{\text{cam}}=6$ surrounding cameras on the autonomous vehicle. For each pair of camera image and LiDAR point cloud, we have the transformation matrix $\mathbf{T}_{n}\in\mathbb{R}^{3\times 4}$ indicating the projection between each camera plane and LiDAR coordinate, where $n=1,2,...,N_{\text{cam}}$.
Encoding. The goal of unsupervised 3D representation learning for fusion perception is to pre-train the LiDAR, camera and fusion encoder in an unsupervised manner. Hence, we first voxelize and embed the raw LiDAR point cloud $\mathbf{P}$ with LiDAR encoder $f^{\text{enc}}_{\text{P}}$
$\hat{\mathbf{P}}=f^{\text{enc}}_{\text{P}}(\mathbf{P}),$ (1)
where $\hat{\mathbf{P}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{P}}}$ is the embedded 3D features for LiDAR point cloud. $\hat{D}$, $\hat{H}$ and $\hat{W}$ are spatial resolutions of the embedded features and $\hat{d}_{\text{P}}$ is number of feature channels after encoding. Then for camera images $\mathcal{I}$, we encode them and project the 2D features to 3D space with $\mathcal{T}$ with image encoder $f^{\text{enc}}_{\text{I}}$
$\hat{\mathbf{I}}=f^{\text{enc}}_{\text{I}}(\mathcal{I},\{\mathbf{T}_{n}\}_{n=1}^{N_{\text{cam}}}),$ (2)
where $\hat{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{I}}}$ is the embedded 3D features for surrounding camera images with the similar dimensions as $\hat{\mathbf{P}}$ except for $\hat{d}_{\text{I}}$ feature channels. With $\hat{\mathbf{P}}$ and $\hat{\mathbf{I}}$, we further concatenate them along feature dimension and apply the fusion encoder $f^{\text{enc}}_{\text{fusion}}$ to get the fusion feature $\hat{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$,
$\hat{\mathbf{F}}=f^{\text{enc}}_{\text{fusion}}([\mathbf{P},\mathbf{I}]).$ (3)
Loss Function. To guide $f^{\text{enc}}_{\text{P}}$, $f^{\text{enc}}_{\text{I}}$ and $f^{\text{enc}}_{\text{fusion}}$ to learn good representations in an unsupervised manner, CLAP first embed the fusion features $\hat{\mathbf{F}}$ with a shallow 3D convolution network $f^{\text{3D}}$ to get $\tilde{\mathbf{F}}=f^{\text{3D}}(\hat{\mathbf{F}})$ and we have $\tilde{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$. Then a rendering loss $\mathcal{L}_{\text{rend}}$ for masked-reconstruction is applied on $\tilde{\mathbf{F}}$. Furthermore, a swapping prototype assignment prediction scheme $\mathcal{L}_{\text{proto}}$ is utilized in order to incorporate interaction of the two modalities into pre-training. The overall loss function is as below:
$\mathcal{L}=\mathcal{L}_{\text{rend}}(\mathbf{P},\tilde{\mathbf{F}},\mathcal{T})+\omega_{\text{proto}}\times\mathcal{L}_{\text{proto}}(\hat{\mathbf{P}},\hat{\mathbf{I}}),$ (4)
with $\omega_{\text{proto}}$ as a weighting parameter to balance the losses.
3.2 Neural Field and Differentiable Rendering
Inspired by the success of UniPAD [53], CLAP applies a differentiable rendering decoder with neural field to conduct the masked-and-reconstruction pre-training scheme. Here we first introduce Neural Field, which is the basis for camera images and point clouds rendering, and then discuss the differentiable rendering process on depth and RGB values.
Neural Field. Given a specific point $\mathbf{p}=[x,y,z]\in\mathbb{R}^{3}$ in the 3D space, the feature $\mathbf{f}_{\text{p}}\in\mathbb{R}^{\hat{d}_{\text{F}}}$ at $\mathbf{p}$ is queried from the fusion 3D embedding $\tilde{\mathbf{F}}$ by trilinear interpolation
$\mathbf{f}_{\text{p}}=f^{\text{tri}}(\mathbf{p},\hat{\mathbf{P}}_{t}),$ (5)
where $f^{\text{tri}}$ is an built-in module implemented in Pytorch [35]. Then we extract geometry features $\mathbf{f}_{\text{geo}}\in\mathbb{R}^{\hat{d}_{\text{geo}}}$ at $\mathbf{p}$ with $f^{\text{geo}}$, taking the concatenation of location $\mathbf{p}$ and queried feature $\mathbf{f}_{\text{p}}$ as inputs,
$\mathbf{f}_{\text{geo}}=f^{\text{geo}}([\mathbf{p},\mathbf{f}]).$ (6)
With geometry features $\mathbf{f}_{\text{geo}}$, we predict the signed distance value $s\in\mathbb{R}$ [4, 28] and color value $s\in\mathbb{R}$ [44] at $\mathbf{p}$ with $f^{\text{SDF}}$ and $f^{\text{RGB}}$. $f^{\text{geo}}$, $f^{\text{RGB}}$ and $f^{\text{SDF}}$ are parameterized by Multi-layer Perceptron,
$s=f^{\text{SDF}}(\mathbf{f}_{\text{geo}})\ \ ,\ \ c=f^{\text{RGB}}(\mathbf{f}_{\text{geo}}).$ (7)
Differentiable Rendering. Similar to [32, 44], we first sample $N_{\text{L}}$ or $N_{\text{C}}$ rays at the LiDAR or camera sensor origin $\mathbf{o}$, each of which is described by its normalized direction $\mathbf{d}$ and $\mathbf{o}$. Next, we sample $N_{\text{ray}}$ points following [44] along each ray. Here each point along the ray can be interpreted by $\mathbf{p}=o+r\mathbf{d}$, where $r$ is the range from the sensor origin to the point $\mathbf{p}$. Thus the sampled point set can be annotated by $\{\mathbf{p}_{n}=\mathbf{o}+r_{n}\mathbf{d}\}_{n=1}^{N_{\text{ray}}}$ and we predict the estimated signed distance value $s_{n}$ and color value $c_{n}$ for them with $f^{\text{geo}}$, $f^{\text{RGB}}$ and $f^{\text{SDF}}$. Following [44], we estimate the occupancy value $\alpha_{n}$ for each sampled point,
$\alpha_{n}=\max{(\frac{\Phi_{h}(s_{n})-\Phi_{h}(s_{n+1})}{\Phi_{h}(s_{n})},0)}.$ (8)
Here $\Phi_{h}(x)=(1+e^{-hx})^{-1}$ stands for the sigmoid function paired with a learnable scalar $h$. After that, we predict the accumulated transmittance $\mathcal{T}_{n}$ similar to [44]
$\mathcal{T}_{n}=\prod^{n-1}_{i=1}(1-\alpha_{i}).$ (9)
Based on $\mathcal{T}_{n}$, we compute an unbiased and occlusion-aware weight $w_{n}=\mathcal{T}_{n}\alpha_{n}$ [44] and integrate all samples along the ray to predict the range $\tilde{r}$ or color $\tilde{c}$ along this ray,
$\tilde{r}=\sum_{n=1}^{N_{\text{ray}}}w_{n}*r_{n}\ \ ,\ \ \tilde{c}=\sum_{n=1}^{N_{\text{ray}}}w_{n}*c_{n}$ (10)
For the observed LiDAR points, it is evident that signed distance value at those points are 0. The final loss function for differentiable rendering is a combination of L-1 loss on range and color predictions and the surface SDF loss.
$\mathcal{L}_{\text{rend}}=\frac{1}{N_{\text{L}}}\sum_{i=1}^{N_{\text{L}}}(\omega_{\text{L}}|r_{i}-\tilde{r}_{i}|+\omega_{\text{sur}}|s_{i}|)+\omega_{\text{C}}\frac{1}{N_{\text{C}}}\sum_{i=1}^{N_{\text{C}}}|c_{i}-\tilde{c}_{i}|,$ (11)
where $\omega_{\text{L}}$, $\omega_{\text{sur}}$ and $\omega_{\text{C}}$ are weighting parameters to balance the contributions from the three losses. $s_{i}$ is the predicted signed distance value at the observed point along the sampled ray.
Figure 3: Inspiration of Curvature Sampling. From the two zoom-in areas, it can be found that areas with low curvature tend to be less informative, like road plane, while those with high curvature might provide more information for pre-training (objects like cars).
3.3 Curvature Sampling
In order to make joint unsupervised representation learning feasible, we have to make $N_{\text{L}}\ll N_{\text{P}}$ and $N_{\text{C}}\ll H\cdot W\cdot N_{\text{cam}}$. Intuitively, uniform sampling can be used. But due to the relatively small sample number compared to the raw inputs ($\sim\frac{1}{100}$), uniform sampling brings little improvement against separate pre-training, which is contradictory to our motivation. Hence, we need to sample more informative part of the scene for $\mathcal{L}_{\text{rend}}$. As shown in Figure 3, we are inspired by the observation that surface with higher curvature (surface of a vehicle) generally contains more information as compared to that with lower curvature (road plane). Therefore, we propose the Curvature Sampling for effective sampling. For each point $\mathbf{p}$ in the LiDAR point cloud $\mathbf{P}$, we first estimate the surface normal by deriving the signed distance function with respect to $\mathbf{p}$
$\mathbf{n}=\frac{\delta f^{\text{SDF}}([\mathbf{p},\mathbf{f}])}{\delta\mathbf{p}},$ (12)
where $\mathbf{n}\in\mathbb{R}^{3}$ is the predicted normal. Then we normalized $\mathbf{n}$ to get the direction of the normal
$\tilde{\mathbf{n}}=\frac{\mathbf{n}}{||\mathbf{n}||_{2}}.$ (13)
Here $||\cdot||_{2}$ is the L-2 norm. Next we estimate the mean curvature $\mathbf{c}\in\mathbb{R}^{3}$ at $\mathbf{p}$ by applying the vector differential operator on $\tilde{\mathbf{n}}$ with respect to $\mathbf{p}$
$\mathbf{c}=\nabla_{\mathbf{p}}\tilde{\mathbf{n}}.$ (14)
For each point $\mathbf{p}_{n}$ in LiDAR point cloud, we compute the average of $\mathbf{c}_{n}$ as the sampling weights $\omega_{n}$ and sample $N_{\text{L}}$ LiDAR point cloud with a Multinomial Sampler implemented in PyTorch [35] for differentiable rendering. For pixels on image plane, we project the LiDAR point cloud back to image planes with $\mathcal{T}$, assign $\omega_{n}$ of each point to the projected pixel and apply a gaussian blur kernel of size $K_{\text{gaus}}$ to densify the weights, with which we sample $N_{\text{C}}$ pixel for $\mathcal{L}_{\text{rend}}$. As the estimation of curvature here is noisy especially during the first few epochs, we set up a warm-up stage with $N_{\text{warmup}}$ epochs with uniform sampling and after $N_{\text{warmup}}$ epochs, Curvature Sampling is utilized.
3.4 Swapping Prototype Assignment Prediction
CLAP aims to utilize the interaction between camera and LiDAR encoders for joint unsupervised 3D representation learning. Meanwhile, we hope to learn a common feature space standing for segments of different objects in an unsupervised manner. Inspired by SwAV [3], we propose to use learnable prototypes to represent parts of the scenes. We first randomly initialize $N_{\text{K}}$ learnable prototypes $\mathbf{K}\in\mathbb{R}^{N_{\text{K}}\times d_{\text{K}}}$, each of which is a $d_{\text{K}}$ vectors. Then we pre-train the backbones to understand the interaction between different modalities with a swapping prototype assignment prediction loss [3]. Furthermore, we introduce an Expectation-Maximization scheme to optimize the learnable prototypes and a gram matrix minimization loss to avoid collapse [3] of the prototype learning.
Swapping Prototype Assignment Prediction. We project the LiDAR embeddings $\hat{\mathbf{P}}$ and camera embeddings $\hat{\mathbf{I}}$ separately with two projection heads $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ to the same dimension as $\mathbf{K}$
$\dot{\mathbf{P}}=f^{\text{proj}}_{\text{P}}(\hat{\mathbf{P}})\ \ ,\ \ \dot{\mathbf{I}}=f^{\text{proj}}_{\text{I}}(\hat{\mathbf{I}}),$ (15)
where $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ are parameterized by Multi-Layer Perceptron and $\dot{\mathbf{P}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{K}}}$, $\dot{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{K}}}$. We denote $N_{\text{3D}}=\hat{D}\times\hat{H}\times\hat{W}$ and then normalize and reshape the projected embeddings into $\dot{\mathbf{P}}\in\mathbb{R}^{N_{\text{3D}}\times\hat{d}_{\text{K}}}$ and $\dot{\mathbf{I}}\in\mathbb{R}^{N_{\text{3D}}\times\hat{d}_{\text{K}}}$. After that, similarity scores $\mathbf{S}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$ between 3D embeddings and prototypes are computed separately for LiDAR and camera branches
$\mathbf{S}_{\text{P}}=\dot{\mathbf{P}}\cdot\mathbf{K}^{\intercal}\ \ ,\ \ \mathbf{S}_{\text{I}}=\dot{\mathbf{I}}\cdot\mathbf{K}^{\intercal}.$ (16)
Similar to SwAV [3], we detach $\mathbf{S}_{\text{P/I}}$ and apply sinkhorn algorithm [9] to make them approach the double stochastic matrix in $N_{\text{sink}}$ iterations. We denote the updated matrix as codes $\mathbf{Q}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$. The swapping prototype assignment loss is formulated with a temperature parameter $\tau$ [3],
$\begin{split}\mathcal{L}_{\text{SwAV}}&=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\mathbf{Q}_{\text{I}}^{n,m}\log\frac{exp(\mathbf{S}_{\text{P}}^{n,m})/\tau}{\sum_{k=1}^{N_{\text{K}}}exp(\mathbf{S}_{\text{P}}^{n,k})/\tau}\\ &+\mathbf{Q}_{\text{P}}^{n,m}\log\frac{exp(\mathbf{S}_{\text{I}}^{n,m})/\tau}{\sum_{k=1}^{N_{\text{K}}}exp(\mathbf{S}_{\text{I}}^{n,k})/\tau}\}.\end{split}$ (17)
Expectation-Maximization. In order to guide the learnable prototypes to represent parts of the environment, we introduce the Expectation-Maximization algorithm [33] to further optimize the prototypes. In the Expectation step, we compute the probability $\hat{\mathbf{S}}_{\text{P/I}}$ that each prototype is assigned to each embeddings by applying a softmax operation on $\mathbf{S}_{\text{P/I}}$. Then for Maximization step, we expect to maximize the probability of the assignment between one prototype to one specific part of the scene and this is the same as minimizing the entropy of the similarity matrix. Hence, the EM loss is computed as,
$\mathcal{L}_{\text{EM}}=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\hat{\mathbf{S}}_{\text{P}}^{n,m}log\hat{\mathbf{S}}_{\text{P}}^{n,m}+\hat{\mathbf{S}}_{\text{I}}^{n,m}log\hat{\mathbf{S}}_{\text{I}}^{n,m}\}$ (18)
Gram Matrix Minimization. When training the randomly initialized prototypes, the network might learn a short cut with all prototypes being the same [3], which is called collapse. To avoid this, we estimate similarity between prototypes by the gram matrix $\mathbf{G}=\mathbf{K}\mathbf{K}^{\intercal}$ of prototypes $\mathbf{K}$, the dimension of which is $\mathbf{G}\in\mathbb{R}^{N_{\text{K}}\times N_{\text{K}}}$. Finally we minimize the average of the non-diagonal elements of $\mathbf{G}$ in order to avoid collapse
$\mathcal{L}_{\text{GMM}}=\frac{1}{N_{\text{K}}(N_{\text{K}}-1)}\sum_{n}^{N_{\text{K}}}\sum_{m=1,m\neq n}^{N_{\text{K}}}\mathbf{G}^{n,m}.$ (19)
Overall Prototype Learning Loss. We apply weighting parameters $\omega_{\text{SwAV}}$, $\omega_{\text{EM}}$ and $\omega_{\text{GMM}}$ to balance the three losses proposed above, which leads to the overall loss function for prototype learning,
$\mathcal{L}_{\text{proto}}=\omega_{\text{SwAV}}\mathcal{L}_{\text{SwAV}}+\omega_{\text{EM}}\mathcal{L}_{\text{EM}}+\omega_{\text{GMM}}\mathcal{L}_{\text{GMM}}.$ (20)
Init. mAP NDS Car Truck C.V. Bus Trailer Barrier Mot. Bic. Ped. T.C.
Rand. 48.69 55.28 78.52 46.64 16.18 50.44 22.11 57.00 46.87 30.56 76.16 62.40
UniPAD 49.81 +0.72 55.29 +0.01 80.81 42.81 17.08 48.98 25.85 61.72 50.19 27.53 78.53 64.57
CLAP 51.17 +2.48 57.04 +1.76 79.56 48.43 18.84 56.34 23.98 60.60 48.87 34.11 78.08 62.87
Table 1: Results for fine-tuning on 5% of training set in NuScenes [2]. “Init.” means the way to initialize models, including the baseline UniPAD [53]. To avoid the case the performance difference is brought by accelerating convergence, we first increase the training iteration of training-from-scratch model (BEVFusion [27]) until we observe convergence, which leads to the “Rand.” row. Then we pre-train the backbones with UniPAD [53] and the proposed CLAP, initialize BEVFusion with the pre-trained weights and then fine-tune the model with the same number iterations as “Rand.”. We provide mAP and NDS as an evaluation of the overall performance of different models and highlight the best mAP and NDS with bold font. We also indicate the performance improvement by green color. For detailed categories, we provide AP for them and “C.V.”, “Mot.”, “Bic.”, “Ped.” and “T.C.” are abbreviations for Construction Vehicle, Motorcycle, Bicycle, Pedestrian and Traffic Cone. All the results here are in %.
4 Experiments
Unsupervised 3D representation learning for fusion perception aims to pre-train both LiDAR and camera encoders and initialize downstream models with the pre-trained weights to gain performance improvement in downstream tasks. In this section, we design extensive experiments on the popular autonomous driving dataset NuScenes [2] to demonstrate the effectiveness of CLAP. To begin with, we describe experiment setups in Section 4.1. Next, we show and analyze main results in Section 4.2. Finally, we provide ablation study and visualizations separately in Section 4.3 and 4.4.
4.1 Settings
Datasets. We use the popular autonomous driving dataset NuScenes [2] to evaluate the performance of CLAP. NuScenes [2] uses one roof LiDAR and six surrounding cameras to collect data. The LiDAR is a 32-beam Velodyne and collecting frequency is 20Hz. The frequency of camera capturing is 12Hz. [2] conducts the synchronization and provides paired data of LiDAR point cloud and camera images. The whole NuScenes dataset contains 1000 scenes collected in Boston and Singapore. Each scene lasts for around 20 seconds and there are a total of 5.5 hours data. Following convention practice from [2, 42], we divide the whole dataset into training set with 850 scenes and validation set with 150 scenes. Pre-trainings are conducted on the whole training set and downstream task is 3D object detection under few-shot setting, where we randomly sample $5\%$ labeled data in the training set to train the randomly initialized model and fine-tune the pre-trained ones. The overall mAP (mean accurate precisions) and NDS (NuScenes Detection Score) are reported, along with AP (accurate precisions) for different categories.
Downstream 3D Object Detectors. We conduct experiments on one of the SOTA fusion 3D object detector called BEVFusion [27]. BEVFusion is implemented in the popular code repository for autonomous driving perception called OpenPCDet [42]. We use average precisions of various categories (APs), mean average precision (mAP) and NuScenes Detection Score (NDS) [2] as the evaluation metrics. We follow a similar setting in [48] to gradually increase training iterations of the model without pre-trained weights until convergence is observed. Here convergence means further increasing training iterations will not improve the performance. Then the number of training iterations is fixed for fine-tuning pre-trained models. This setting avoid the case that pre-training only accelerate convergence and make sure that pre-training indeed improve the performance of downstream models, that is improving the sample efficiency of the downstream task.
Baseline Pre-training Method for Fusion Perception. We incorporate UniPAD [53] as the baseline 3D unsupervised pre-training method. We utilize the official implementation from [53] to pre-train the backbones of BEVFusion [27]. As pre-trained image backbones are broadly used for image feature extraction, the implementation of [53] also use a pre-trained image backbone from [8] to initiate their pre-training and the training set for this backbone does not involve any data from NuScenes [2]. We adopt this practice and also for training-from-scratch model, we initialize the camera encoder with the same pre-trained weights from [8].
Implementation Details of CLAP. The feature channels for LiDAR embeddings $\hat{\mathbf{P}}$, image embeddings $\hat{\mathbf{I}}$, fusion embeddings $\hat{\mathbf{F}}$ and swapping prototype assignment prediction are respectively set to $\hat{d_{\text{P}}}=256$, $\hat{d_{\text{I}}}=80$, $\hat{d_{\text{F}}}=512$ and $\hat{d}_{\text{K}}=128$. Sampling number for LiDAR point cloud and camera pixel are $N_{\text{L}}=8192$ and $N_{\text{C}}=1024\times N_{\text{cam}}$, where $N_{\text{cam}}=6$ for NuScenes dataset [2]. The number of sample points along each ray is $N_{\text{ray}}=96$. Warm-up epochs for Curvature Sampling is $N_{\text{warmup}}=4$. For swapping prototype assignment prediction, we set the number of learnable prototypes and sinkhorn update iterations to $N_{\text{K}}=512$ and $N_{\text{sink}}=3$. The temperature for swapping prediction loss is $\tau=1.0$. The loss weighting parameters are implemented as $\omega_{\text{proto}}=1.0$, $\omega_{\text{L}}=2.0$, $\omega_{\text{sur}}=0.1$, $\omega_{\text{C}}=0.1$, $\omega_{\text{SwAV}}=1.0$, $\omega_{\text{EM}}=0.1$ and $\omega_{\text{GMM}}=0.1$. We use a learning rate of 0.00005 with a cosine learning schedule for pre-training and use mask augmentation for CLAP with a masking rate of 0.9.
4.2 Main Results
Both CLAP and baseline methods are pre-trained on the whole training set of NuScenes [2] without labels. To compare the effectiveness of CLAP and UniPAD [53], we randomly sample 5% of the training labels and fine-tune BEVFusion [27] initialized by the pre-trained weights. Furthermore, as the size of pre-training dataset is limited, we randomly sample 2.5%, 1% and 0.5% of the training labels and fine-tune BEVFusion with CLAP pre-trained weights to see whether potential scaling property exists in pre-training with CLAP.
Comparison to Previous SOTA Method. As shown in Table 1, CLAP achieves 2.48% mAP improvement over randomly initialization at convergence, which is $300\%$ more improvement for mAP than SOTA unsupervised 3D representation method UniPAD [53]. For NDS metric, UniPAD [53] only achieves comparable performance while CLAP surpasses the train-from-scratch model by 1.76% . When it turns to different categories, CLAP generally benefit the performance of all the categories and for Construction Vehicle, Bus, Barrier, Motorcycle and Bicycle, the improvement over random initialization are more than 2% AP.
Potential Scaling Property. As we are not able to scale up the pre-training dataset at current stage, we explore potential scaling property by gradually decreasing the sample numbers (2.5%, 1% and 0.5%) for fine-tuning, which increases the ratio between pre-training data and fine-tuning data. The results are shown in Table 2 together with the results of fine-tuning on 5% of training set in NuScenes. It can be found that as the ratio between pre-training data and fine-tuning data gets larger, the performance improvement by CLAP increases and CLAP provides a gain up to 7.22% mAP and 4.71% NDS with 0.5% fine-tuning data. These results show that CLAP is promising in scaling property and in the future, if we can scale up the pre-training dataset, CLAP might further improve current SOTA performance.
Init. Fine-tune Data mAP NDS
Random 5% 48.69 55.28
CLAP 51.17 +2.48 57.04 +1.76
Random 2.5% 39.12 40.01
CLAP 42.86 +3.74 42.18 +2.17
Random 1% 26.22 29.82
CLAP 30.53 +4.31 31.87 +2.05
Random 0.5% 16.49 22.61
CLAP 23.71 +7.22 27.32 +4.71
Table 2: Results on potential scaling property of CLAP . “Init.” means the method for initialization. “Fine-tune Data” stands for the sample number of training set in NuScenes [2] used for fine-tuning/training-from-scratch. We gradually increase the number of training iterations for training-from-scratch model (BEVFusion [27]) with different numbers of training data until we observe convergence, leading to the results of “Rand.”. Then we initialize BEVFusion with weights pre-trained by CLAP and then fine-tune the model with the same numbers of iterations as “Rand.”. We provide mAP and NDS as an evaluation of the overall performance. All the results here are in %.
Figure 4: Visualization of curvature estimation. The color of the points change from blue to red, indicating lower sample weight / curvature to higher ones. We highlight those correct regions with orange boxes and those noisy estimation with green ones. Some of the correct estimated regions are further zoom-in for better understanding. Best view in color.
4.3 Visualizations
In this section, we provide visualization to show the effectiveness of the proposed Curvature Sampling. In order to evaluate whether the estimated curvature is able to help sampling more informative part of the environment, we use the pre-trained model to estimate the curvature of LiDAR point clouds and use heatmap color to indicate the weight computed in Section 3.3. The visualization results are shown in Figure 4. We use orange boxes to highlight those regions with relatively correct estimation and green ones for those with noisy estimation. Furthermore, for better understanding about the curvature estimation, we zoom in for some correctly estimated regions. It can be found that though some noise exists, CLAP is able to predict high weights for those highly informative region for sampling and meanwhile assign lower weights to most of the background, which makes joint pre-training feasible.
Joint Pre-train Cur. Sam. Proto. Learning mAP
49.81
49.55
50.81
51.17
[…]
4.4 Ablation Study
We conduct ablation study to evaluate the effectiveness of different components. As shown in Table 3, it can be found that using uniform sampling (second line) to make joint pre-training feasible does not bring improvement over separate pre-training (first line) [53]. Then, Curvature Sampling (third line) improves the performance over uniform sampling and separate pre-training by assigning larger weights to more informative parts of the 3D scene during sampling. Finally, the Prototype Learning scheme (fourth line) introduces interaction of LiDAR and camera encoders into pre-training and achieve the best performance.
5 Conclusion
In this paper, we propose CLAP to jointly pre-train for fusion perception in an unsupervised manner via differentiable rendering. To make joint pre-training feasible, CLAP utilizes the novel Curvature Sampling to sample more informative parts of the 3D scene for pre-training. Further, a swapping prototype prediction scheme is introduced in order to make full use of the interaction between LiDAR point clouds and camera images. To make the learnable prototypes better represent parts/semantics of objects, CLAP brings the idea of Expectation-Maximization for prototypes optimization, alongside with a gram matrix regularization term to avoid collapse. Experiment results demonstrate that CLAP is superior in unsupervised 3D representation learning and has the potential to scale up. In the future, when a larger pre-training dataset is available, we expect to scale up the pre-training with CLAP for more performance improvement.
Supplementary Material
Init. F.T. mAP Vehicle Pedestrian Cyclist
0-30m 30-50m 50m- 0-30m 30-50m 50m- 0-30m 30-50m 50m-
Rand* 5% 20.48 58.03 25.22 12.98 11.62 9.75 6.97 21.55 6.83 3.11
CLAP* 22.86 +2.38 58.37 26.38 14.07 12.60 9.50 7.88 30.08 10.50 5.39
Rand 46.07 76.71 51.15 31.84 37.53 20.12 9.84 62.00 42.61 24.18
CLAP 46.88 +0.81 76.98 51.64 31.31 38.79 20.60 9.74 63.75 43.21 26.83
Rand* 100% 64.00 86.21 70.20 58.20 57.80 41.18 23.55 75.95 61.45 45.80
CLAP* 64.74 +0.74 88.14 72.59 59.13 57.37 42.24 24.22 77.11 61.91 45.63
Rand 65.03 88.18 74.23 61.75 57.32 38.90 21.96 78.07 64.32 48.16
CLAP 65.56 +0.53 87.97 72.77 62.11 58.33 40.11 21.29 78.63 64.70 47.27
Table 4: Results for transferring experiments on Once [29] dataset. CenterPoint [54] is used as the downstream detector. “Init.” indicates the initialization methodshorthands. “F.T.” indicates the number of training samples in fine-tuning stage. Overall mAP and APs for different categories within different ranges are shown in this table. “Rand*” means training the randomly initialized model with the original training iterations in OpenPCDet [42]. “Rand” indicates that we increase the number of training iterations for randomly initialized model until convergence is observed. “CLAP*” indicates that we pre-train the backbones with CLAP on NuScenes [2] and then fine-tune on Once with the original iterations in [42]. “CLAP” uses the same fine-tuning iterations as “Rand”. We use green color to highlight the performance improvement brought by CLAP. All the results are in %.
A Transferring to other datasets
We conduct further experiments to evaluate the transferring ability of CLAP. Specifically, we select LiDAR-based 3D object detector CenterPoint [54] on Once [29] for the downstream task. A 40-beam LiDAR is utilized in Once [29] to collect 15k labeled training data. We randomly sample $5\%$ and also use all of the labeled training set to train the from-scratch model until convergence is observed. Then we use pre-trained weights by CLAP on NuScenes [2] to initialize the same model and fine-tune it with the same training iterations as the randomly initialized model. Results are shown in Table 4. It can be found that pre-training by CLAP also benefits LiDAR-based 3D object detection, even in a cross-dataset setting. And if we look at the performance of “Rand*” and “CLAP*”, CLAP also accelerates the convergence in downstream task.
B Visualization of Prototype Learning
We use the model pre-trained by CLAP to infer the 3D features and assign prototypes to different LiDAR points in the 3D space. Then we use different colors to indicate different prototypes and visualize them in Figure 5. If can be found that the background road plane inside the same frame is generally assigned to the same prototype. And foreground vehicles are assigned to another prototype. This demonstrate that the proposed prototype learning scheme actually learns some level of semantics understanding in an unsupervised manner. However, as our pre-training does not incorporate any label, it can also be found that the prototype assignment has some noise, for example some of the road plane points are assigned to other prototypes.
Figure 5: Visualization of the prototype learning results. Different color indicates different prototype assignments.

### v2

@@
 Abstract
-Unsupervised 3D representation learning via masked-and-reconstruction with differentiable rendering is promising to reduce the labeling burden for fusion 3D perception.
-However, previous literature conduct pre-training for different modalities separately because of the hight GPU memory consumption.
-Consequently, the interaction between the two modalities (images and point clouds) is neglected during pre-training.
-In this paper, we explore joint unsupervised pre-training for fusion 3D perception via differentiable rendering and propose CLAP, short for Curvature sampLing and swApping Prototype assignment prediction.
-The contributions are three-fold.
-1) To overcome the GPU memory consumption problem, we propose Curvature Sampling to sample the more informative points/pixels for pre-training.
-2) We propose to use learnable prototypes to represent parts of the scenes in a common feature space and bring the idea of swapping prototype assignment prediction to learn the interaction between the two modalities.
-3) To further optimize learnable prototypes, we propose an Expectation-Maximization training scheme to maximize the similarity between embeddings and prototypes, followed by a Gram Matrix Regularization Loss to avoid collapse.
-Experiment results on NuScenes show that CLAP achieves $300\%$ more performance gain as compared to previous SOTA 3D pre-training method via differentiable rendering.
+Unsupervised 3D representation learning reduces the burden of labeling multimodal 3D data for fusion perception tasks.
+Among different pre-training paradigms, differentiable-rendering-based methods have shown most promise.
+However, existing works separately conduct pre-training for each modalities due to computational costs of processing large point clouds with images.
+As such, mutual benefit of high-level semantics (from image) and 3D structure (from point cloud) has not been exploited.
+To address this gap, we propose a joint unsupervised differentiable-rendering-based pre-training method for images and point clouds, termed CLAP, short for Curvature sampLing and leArnable Prototype.
+Specifically, our method overcomes the computational hurdle by Curvature Sampling to select the more informative points/pixels for pre-training.
+To uncover the performance benefits brought by their complementarity, we propose to use learnable prototypes to represent parts of the 3D scenes in a common feature space and an Expectation-Maximization training scheme to associate embeddings of each modality to prototypes.
+We further propose a swapping prediction loss that explores their interplay through prototypes along with a Gram Matrix Regularization term to maintain training stability.
+Experiments on NuScenes and Waymo datasets show that CLAP achieves up to $100\%$ more performance gain as compared to previous SOTA pre-training methods.
 Codes and models will be released.
 1 Introduction
-3D perception provides understanding of the ego vehicle’s surrounding 3D space and plays an important role in downstream prediction and control tasks in autonomous driving.
-Single-modality 3D perception has long been studied with camera [11, 6, 40, 45, 55, 19] or LiDAR (Light-Detection-And-Ranging) [36, 50, 39, 54, 12, 1] inputs, which is less robust to changing lighting and weather conditions [30].
-To compensate shortages of different modalities, research prevails on fusion 3D perception with camera images and LiDAR point clouds [27, 25, 21, 24, 20, 31, 7].
-Figure 1: Unlike previous unsupervised 3D representation learning method UniPAD [53] that separately pre-train LiDAR and camera encoders (a), our proposed method CLAP conducts joint pre-training for fusion perception.
-Meanwhile, labeling in 3D space is notoriously time-and-energy-consuming.
-Unsupervised 3D representation learning [23, 18, 5, 52, 48, 17, 57], which pre-trains backbones without any label and applies the pre-trained weight to initialize downstream models for performance improvement, shows potential to alleviate labeling burden in 3D perception.
-Hence, the community starts to explore pre-training for fusion 3D perception in an unsupervised manner [41, 56, 22, 53], among which UniPAD [53] shows superior performance.
-UniPAD [53] is a masked-reconstruction-based method and utilizes differentiable rendering as decoder during pre-training.
-However, during pre-training stage, the camera and LiDAR encoders are separately pre-trained without considering the interaction between the two modalities in [53].
-The reason stems from the high GPU memory consumption.
-If all of the LiDAR points and camera pixels are considered for pre-training, the most advanced GPU up-to-date (H100) is only able to hold batchsize=1.
-But batchsize is an important hyperparameter for image encoder and a small one normally brings degradation to the performance, making it unfeasible for 3D pre-training.
-In this paper, we explore joint unsupervised pre-training for fusion 3D perception via differentiable rendering and propose CLAP, which is shorthand for Curvature sampLing and swApping Prototype assignment prediction.
-To begin with, we need to sample the LiDAR points and camera pixels for pre-training in order to make joint pre-training feasible.
-The most direct way is to use uniform sampling but as the sample number is small compared to the raw inputs ($\sim\frac{1}{100}$), uniform sampling does not brings improvement against separate pre-training.
-In order to sample more informative parts of the environment for pre-training, CLAP first estimates curvature of each LiDAR point in the 3D space by taking second order derivative of the SDF (signed distance field) function.
-Then we compute weights proportional to the estimated curvature and sample more informative points according to these weights.
-As for pixel sampling, we project the LiDAR points onto image planes and assign the weights to pixel.
-As LiDAR point clouds are sparse, we further apply gaussian blur to broadcast the weights to nearby pixels and sample pixels according to these weights.
-With Curvature Sampling, we further consider how to better utilize the interaction between camera images and LiDAR point clouds to gain some level of understanding about “objectness” or “parts of objects” in an unsupervised way.
-Inspired by SwAV [3], we propose to use learnable prototypes to learn to represent parts of the scenes.
-Similar to [3], CLAP applies a swapping prototype assignment prediction loss between camera and LiDAR embeddings to learn the common feature space.
-To further optimize the learnable prototypes, an Expectation-Maximization training scheme is proposed.
-In the Expectation step, we compute the probability that each prototype is assigned to the 3D embeddings in the 3D space.
-Then in the Maximization step, we maximize the similarity between embeddings and prototypes by minimizing the entropy of the assignment scores.
-Finally, to avoid collapse of the prototype [3], a gram matrix regularization Loss is proposed to minimize the correspondence among prototypes.
-Through extensive experiments on the popular autonomous driving dataset NuScenes [2], we demonstrate that CLAP is able to achieve up to $300\%$ more improvement than previous SOTA method [53].
+3D perception facilitates spatial applications such as autonomous driving.
+The autonomous system, on which these applications are deployed, are typically equipped with multiple sensors including visual ones like cameras that produce RGB images, and range sensors like LiDAR (Light-Detection-And-Ranging) that generate point clouds.
+Fusion of these two modalities [30, 26, 22, 25, 21, 32, 8] have generally improved over use of a single modality, e.g., camera [11, 7, 43, 50, 60, 20] or LiDAR [39, 55, 42, 59, 12, 1] separately, in 3D perception performance.
+Figure 1: Unlike previous SOTA unsupervised 3D representation learning method UniPAD [58] that separately pre-train LiDAR and camera encoders with differential rendering (a), our proposed method CLAP conducts joint pre-training for fusion perception.
+However, training multimodal 3D perception models is expensive as labeling in 3D space is notoriously time-and-energy-consuming.
+Unsupervised 3D representation learning, which pre-trains backbones without any label and fine-tuned the pre-trained weight for downstream performance improvement, has shown the potential to alleviate the labeling burden in 3D perception.
+Amongst the many unsupervised 3D pre-training methods [24, 19, 6, 57, 53, 18, 62, 58], use of mask auto-encoding (reconstruction) and differntiable rendering has emerged as the most performant.
+However, encoding and processing high-dimensional multi-modal data (images and point clouds) is computationally expensive: If one were to pre-train using all points and pixels within the point cloud and image, respectively, even the most advanced GPU to date (H100) is only able to hold a batch size of 1.
+Therefore, existing methods, e.g., UniPad [58], has conventionally pre-trained each modality separately.
+While the two modalities are complementary (semantics from image, and structure from point clouds), current methods are unable to exploit their mutual benefits during pre-training.
+To address this pain point, we propose a joint unsupervised pre-training method of image and point cloud modalities based on differentiable rendering.
+Our method, CLAP, short for Curvature sampLing and leArnable Prototype, addresses the primary computational challenge by a curvature sampling strategy, which comes from the observation that there exists redundancy in information when sampling multiple points on the same flat surface.
+This is enabled by estimating the curvature of each point in the 3D space by taking second order derivative of the SDF (signed distance field) function.
+This allows us to identify points at which large change in the surface occurs.
+Through calibration, we can obtain the corresponding pixel within the image, around which we sample assuming the coordinates are normally distributed.
+This sampling strategy captures variations in the point cloud and in turn provide more informative points than the conventional sampling strategy in [58].
+As our curvature sampling strategy has reduced the computational burden to allow both modalities to be simultaneously processed, we propose to model the interplay between image and point cloud modalities through a set of learnable prototypes, which represents parts of the 3D scene and enables a common feature space to bridge the two modalities.
+These prototypes are trained via an Expectation-Maximization (EM) algorithm that maximizes similarity between embeddings for each modality and the set of prototypes.
+Furthermore, we propose to use the swapping prediction loss to explore the interaction between the two modalities.
+Last but not least, we utilize a Gram Matrix Regularization term to minimize similarity across prototypes, which addresses collapse of prototype when trained naively.
+Our contributions are 4-folds: (1) We propose a curvature sampling strategy to identify informative points (and pixels) for sampling, which enables the first joint differentiable-rendering-based pre-training method for fusion perception.
+(2) Learnable prototypes are utilized to learn a common feature space and an Expectation-Maximization approach is proposed to train the prototypes to represent parts of the 3D scene.
+(3) We further propose to use a swapping prediction loss for modality interplay and a Gram Matrix Regularization loss that avoids collapse of prototype learning.
+(4) Through extensive experiments on the popular autonomous driving datasets NuScenes [3] and Waymo [45], we demonstrate the effectiveness of CLAP .
+For example, CLAP achieves up to $100\%$ more improvement than previous SOTA 3D pre-training methods and shows potential scaling property.
 2 Related Work
@@
 For LiDAR-based 3D object detection, there are three main streams with different embedding schemes for point clouds inputs.
-1) Embraced by [36, 38], point-based methods utilize point-level embeddings for 3D object detection.
+1) Point-based methods [39, 41] utilize point-level embeddings for 3D object detection.
 2) Voxel-based methods [55, 10, 59, 1, 56, 12] voxelize the 3D scene and use sparse convolution or transformer for embedding.
@@
 For camera-based 3D perception, [11, 7, 43, 50, 60, 20] embed 2D features on image plane and project these 2D features into 3D space with estimated depth.
-As single-modality perception can be degraded by various weather and lighting conditions [30], fusion 3D perception [7, 27, 25, 21, 24, 20, 31] starts to flourish.
-These methods mainly focus on the way to integrate embeddings of different modalities and use supervised pre-trained 2D and 3D backbones to achieve the SOTA performance.
+Recently, research starts to focus on fusion 3D perception [8, 30, 26, 22, 25, 21, 32] with camera images and LiDAR point clouds as inputs.
+These methods mainly focus on integrating embeddings of different modalities and train both the 2D and 3D backbones in an supervised manner.
 As labeling in 3D space is costly, we explore unsupervised 3D representation learning for fusion 3D perception.
 3D Pre-training.
-Annotating for 3D data is notoriously time-and-energy-consuming and the emergence of unsupervised representation learning for 2D image [14, 43, 3, 13, 46, 15] provides a promising way to alleviate the annotation burden.
-Researcher starts to introduce unsupervised 3D representation learning into scene-level 3D point clouds [47, 26, 16, 5, 18, 23, 34, 52, 48, 17, 57], which can be divided into two contrastive-based and masked-and-reconstruction-based branches.
-Embraced by [47, 26, 16, 5, 18, 23, 34], contrastive-based works propose various ways to build suitable views and conduct contrastive learning to improve the performance in downstream perception task.
+Annotating for 3D data is notoriously time- and energy-consuming and the emergence of unsupervised representation learning for 2D image [15, 47, 4, 13, 51, 16] provides a promising way to alleviate the annotation burden.
+Existing works [52, 27, 17, 6, 19, 24, 28, 38, 36, 57, 53, 18, 62] in unsupervised 3D representation learning into scene-level 3D point clouds can be divided into two contrastive-based and masked-and-reconstruction-based paradigms.
+Contrastive-based works [52, 27, 17, 6, 19, 24, 36, 28] propose various ways to build suitable views and conduct contrastive learning to improve the performance in downstream perception task.
 Inspired by [16] in image domain, [57, 53, 18, 62] propose to first mask the input point clouds and pre-train the 3D encoders with a shallow decoder for reconstructing the unmasked inputs.
@@
 However, due to the high GPU memory consumption, UniPAD [58] is only able to separately pre-train the image and point cloud encoders and fails to utilize the interaction between modalities during pre-training.
-In this paper, we explore unsupervised joint pre-training for 2D and 3D backbones via differentiable rendering.
+In this paper, we explore unsupervised joint pre-training for 2D and 3D backbones via differentiable rendering with Curvature Sampling and Prototype Learning.
+Previous work like SwAV [4] is related to our prototype learning part.
+SwAV [4] explores learnable prototypes to represent different categories in the same modality (image) and uses swapping assignment prediction loss for different views of the same image instance.
+On the contrary, learnable prototypes in CLAP are used to represent part of the 3D scenes and learn the interaction between modalities, which differs from the context in [4].
+Thus we propose a Expectation-Maximization training scheme to maximize similarity between prototypes and 3D embeddings.
+Then a swapping prediction loss is used to learn the modality interaction.
+Furthermore, to avoid prototypes collapsing to the same vector, we propose a Gram Matrix Regularization loss.
 3 Method
@@
 To make joint pre-training feasible, we propose Curvature Sampling to sample informative parts of the 3D scene, as descriped in (a).
-Furthermore, we propose to use learnable prototypes to represent parts of objects and use a swapping prototype assignment prediction loss to incorporate the interaction of different modalities into pre-training.
-Finally, we optimize the learnable prototypes with an Expectation-Maximization scheme together with a gram matrix regularization loss.
-In this section, we introduce CLAP for unsupervised 3D pre-training via differentiable rendering on fusion 3D perception.
+Furthermore, we propose to use learnable prototypes to represent parts of objects in a common feature space and utilize an Expectation-Maximization approach to maximize the similarity between prototypes and 3D embeddings of each modality.
+To delve deeper into the interplay of image semantics and LiDAR geometry, we use swapping prototype prediction loss.
+Finally, we propose a Gram Matrix Regularization loss to prevent collapse of prototype learning.
+In this section, we introduce CLAP for joint unsupervised 3D pre-training via differentiable rendering on fusion 3D perception.
 As described in Fig.
 2, CLAP pre-trains the image, LiDAR and fusion encoders jointly with Neural Field Rendering.
-In order to make joint pre-training feasible, Curvature Sampling is proposed as shown in Fig.
-2 (a).
-To further make use of both modalities, we propose a swapping prototype assignment prediction scheme, which is visualized in Fig.
+In order to enable joint pre-training, Curvature Sampling is proposed as shown in (a) of Fig.
+2.
+To further make use of both modalities, we utilize learnable prototypes to represent parts of the 3D scenes as shown in Fig.
 2 (b).
+To optimize the learnable prototype, we train a common feature space with an Expectation-Maximization training scheme and incorporate interaction between modalities by a swapping prediction loss.
+Finally a Gram Matrix Regularization loss is proposed to avoid collapse in prototype learning.
 We first discuss the formulation and overall pipeline in Section 3.1.
 Then we introduce the details about neural field and differentiable rendering in Section 3.2.
-Finally, we describe the curvature sampling and swapping prototype assignment prediction separately in Section 3.3 and 3.4.
+Finally, we describe the curvature sampling and prototype learning separately in Section 3.3 and 3.4.
 3.1 Formulation and Pipeline
@@
 For example, in NuScenes [3] dataset, $d=2$ represents the intensity and timestamp of each point and there are $N_{\text{cam}}=6$ surrounding cameras on the autonomous vehicle.
-For each pair of camera image and LiDAR point cloud, we have the transformation matrix $\mathbf{T}_{n}\in\mathbb{R}^{3\times 4}$ indicating the projection between each camera plane and LiDAR coordinate, where $n=1,2,...,N_{\text{cam}}$.
+For each pair of camera image and LiDAR point cloud, we have the transformation matrix $\mathbf{T}_{n}\in\mathbb{R}^{3\times 4}$ indicating the projection between the camera plane and LiDAR coordinate, where $n=1,2,...,N_{\text{cam}}$.
 Encoding.
@@
 $\hat{D}$, $\hat{H}$ and $\hat{W}$ are spatial resolutions of the embedded features and $\hat{d}_{\text{P}}$ is number of feature channels after encoding.
-Then for camera images $\mathcal{I}$, we encode them and project the 2D features to 3D space with $\mathcal{T}$ with image encoder $f^{\text{enc}}_{\text{I}}$
-$\hat{\mathbf{I}}=f^{\text{enc}}_{\text{I}}(\mathcal{I},\{\mathbf{T}_{n}\}_{n=1}^{N_{\text{cam}}}),$ (2)
+Then for camera images $\mathcal{I}$, $f^{\text{enc}}_{\text{I}}$ encodes them with swin transformer [29] and uses $\mathcal{T}=\{\mathbf{T}_{n}\}_{n=1}^{N_{\text{cam}}}$ to project the 2D features to 3D space.
+$\hat{\mathbf{I}}=f^{\text{enc}}_{\text{I}}(\mathcal{I},\mathcal{T},\mathbf{P}),$ (2)
 where $\hat{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{I}}}$ is the embedded 3D features for surrounding camera images with the similar dimensions as $\hat{\mathbf{P}}$ except for $\hat{d}_{\text{I}}$ feature channels.
-With $\hat{\mathbf{P}}$ and $\hat{\mathbf{I}}$, we further concatenate them along feature dimension and apply the fusion encoder $f^{\text{enc}}_{\text{fusion}}$ to get the fusion feature $\hat{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$,
+The projection of 2D features to 3D space is similar to [30]: we transform LiDAR points back to image planes with $\mathcal{T}$ and use the projected depths to project the 2D features to 3D space.
+With $\hat{\mathbf{P}}$ and $\hat{\mathbf{I}}$, we further concatenate them along feature dimension and apply the fusion encoder $f^{\text{enc}}_{\text{fusion}}$ to get the fusion feature $\hat{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$ with $\hat{d}_{\text{F}}$ feature dimensions,
 $\hat{\mathbf{F}}=f^{\text{enc}}_{\text{fusion}}([\mathbf{P},\mathbf{I}]).$ (3)
@@
 To guide $f^{\text{enc}}_{\text{P}}$, $f^{\text{enc}}_{\text{I}}$ and $f^{\text{enc}}_{\text{fusion}}$ to learn good representations in an unsupervised manner, CLAP first embed the fusion features $\hat{\mathbf{F}}$ with a shallow 3D convolution network $f^{\text{3D}}$ to get $\tilde{\mathbf{F}}=f^{\text{3D}}(\hat{\mathbf{F}})$ and we have $\tilde{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$.
-Then a rendering loss $\mathcal{L}_{\text{rend}}$ for masked-reconstruction is applied on $\tilde{\mathbf{F}}$.
-Furthermore, a swapping prototype assignment prediction scheme $\mathcal{L}_{\text{proto}}$ is utilized in order to incorporate interaction of the two modalities into pre-training.
+Then a rendering loss $\mathcal{L}_{\text{rend}}$ is applied on $\tilde{\mathbf{F}}$ for masked-reconstruction on both point clouds $\mathbf{P}$ and images $\mathcal{I}$.
+Furthermore, a prototype learning scheme $\mathcal{L}_{\text{proto}}$ is utilized in order to bridge the two modalities and incorporate interaction of image semantics and LiDAR geometry into pre-training.
 The overall loss function is as below:
-$\mathcal{L}=\mathcal{L}_{\text{rend}}(\mathbf{P},\tilde{\mathbf{F}},\mathcal{T})+\omega_{\text{proto}}\times\mathcal{L}_{\text{proto}}(\hat{\mathbf{P}},\hat{\mathbf{I}}),$ (4)
-with $\omega_{\text{proto}}$ as a weighting parameter to balance the losses.
+$\mathcal{L}=\omega_{\text{r}}\times\mathcal{L}_{\text{rend}}(\mathbf{P},\tilde{\mathbf{F}},\mathcal{I})+\omega_{\text{proto}}\times\mathcal{L}_{\text{proto}}(\hat{\mathbf{P}},\hat{\mathbf{I}}),$ (4)
+with $\omega_{\text{r}}$ and $\omega_{\text{proto}}$ as weighting parameters to balance the losses.
 3.2 Neural Field and Differentiable Rendering
-Inspired by the success of UniPAD [53], CLAP applies a differentiable rendering decoder with neural field to conduct the masked-and-reconstruction pre-training scheme.
+Inspired by the success of UniPAD [58], CLAP applies a differentiable rendering decoder with neural field to conduct the masked-and-reconstruction pre-training.
+Different from [58], an additional surface signed distance field loss is applied to better optimize the scene geometry.
 Here we first introduce Neural Field, which is the basis for camera images and point clouds rendering, and then discuss the differentiable rendering process on depth and RGB values.
@@
 where $f^{\text{tri}}$ is an built-in module implemented in Pytorch [37].
-Then we extract geometry features $\mathbf{f}_{\text{geo}}\in\mathbb{R}^{\hat{d}_{\text{geo}}}$ at $\mathbf{p}$ with $f^{\text{geo}}$, taking the concatenation of location $\mathbf{p}$ and queried feature $\mathbf{f}_{\text{p}}$ as inputs,
-$\mathbf{f}_{\text{geo}}=f^{\text{geo}}([\mathbf{p},\mathbf{f}]).$ (6)
-With geometry features $\mathbf{f}_{\text{geo}}$, we predict the signed distance value $s\in\mathbb{R}$ [4, 28] and color value $s\in\mathbb{R}$ [44] at $\mathbf{p}$ with $f^{\text{SDF}}$ and $f^{\text{RGB}}$.
-$f^{\text{geo}}$, $f^{\text{RGB}}$ and $f^{\text{SDF}}$ are parameterized by Multi-layer Perceptron,
-$s=f^{\text{SDF}}(\mathbf{f}_{\text{geo}})\ \ ,\ \ c=f^{\text{RGB}}(\mathbf{f}_{\text{geo}}).$ (7)
+Taking the concatenation of location $\mathbf{p}$ and queried feature $\mathbf{f}_{\text{p}}$ as inputs, we predict the signed distance value $s\in\mathbb{R}$ [5, 31] and color value $c\in\mathbb{R}^{3}$ [49] at $\mathbf{p}$ with $f^{\text{SDF}}$ and $f^{\text{RGB}}$.
+$s=f^{\text{SDF}}([\mathbf{p},\mathbf{f}])\ \ ,\ \ c=f^{\text{RGB}}([\mathbf{p},\mathbf{f}]),$ (6)
+where $f^{\text{RGB}}$ and $f^{\text{SDF}}$ are parameterized by Multi-layer Perceptron.
 Differentiable Rendering.
@@
 Next, we sample $N_{\text{ray}}$ points following [49] along each ray.
-Here each point along the ray can be interpreted by $\mathbf{p}=o+r\mathbf{d}$, where $r$ is the range from the sensor origin to the point $\mathbf{p}$.
-Thus the sampled point set can be annotated by $\{\mathbf{p}_{n}=\mathbf{o}+r_{n}\mathbf{d}\}_{n=1}^{N_{\text{ray}}}$ and we predict the estimated signed distance value $s_{n}$ and color value $c_{n}$ for them with $f^{\text{geo}}$, $f^{\text{RGB}}$ and $f^{\text{SDF}}$.
+Here each point along the ray can be interpreted by $\mathbf{p}=\mathbf{o}+r\mathbf{d}$, where $r$ is the range from the sensor origin to the point $\mathbf{p}$.
+Thus the sampled point set can be annotated by $\{\mathbf{p}_{n}=\mathbf{o}+r_{n}\mathbf{d}\}_{n=1}^{N_{\text{ray}}}$ and we predict the estimated signed distance value $s_{n}$ and color value $c_{n}$ for them with $f^{\text{RGB}}$ and $f^{\text{SDF}}$.
 Following [49], we estimate the occupancy value $\alpha_{n}$ for each sampled point,
 $\alpha_{n}=\max{(\frac{\Phi_{h}(s_{n})-\Phi_{h}(s_{n+1})}{\Phi_{h}(s_{n})},0)}.$ (7)
-Here $\Phi_{h}(x)=(1+e^{-hx})^{-1}$ stands for the sigmoid function paired with a learnable scalar $h$. After that, we predict the accumulated transmittance $\mathcal{T}_{n}$ similar to [44]
-$\mathcal{T}_{n}=\prod^{n-1}_{i=1}(1-\alpha_{i}).$ (9)
-Based on $\mathcal{T}_{n}$, we compute an unbiased and occlusion-aware weight $w_{n}=\mathcal{T}_{n}\alpha_{n}$ [44] and integrate all samples along the ray to predict the range $\tilde{r}$ or color $\tilde{c}$ along this ray,
+Here $\Phi_{h}(x)=(1+e^{-hx})^{-1}$ stands for the sigmoid function paired with a learnable scalar $h$. After that, we predict the accumulated transmittance $t_{n}$ similar to [49]
+$t_{n}=\prod^{n-1}_{i=1}(1-\alpha_{i}).$ (8)
+Based on $t_{n}$, we compute an unbiased and occlusion-aware weight $w_{n}=t_{n}\alpha_{n}$ [49] and integrate all samples along the ray to predict the range $\tilde{r}$ or color $\tilde{c}$ along this ray,
 $\tilde{r}=\sum_{n=1}^{N_{\text{ray}}}w_{n}*r_{n}\ \ ,\ \ \tilde{c}=\sum_{n=1}^{N_{\text{ray}}}w_{n}*c_{n}$ (9)
@@
 The final loss function for differentiable rendering is a combination of L-1 loss on range and color predictions and the surface SDF loss.
-$\mathcal{L}_{\text{rend}}=\frac{1}{N_{\text{L}}}\sum_{i=1}^{N_{\text{L}}}(\omega_{\text{L}}|r_{i}-\tilde{r}_{i}|+\omega_{\text{sur}}|s_{i}|)+\omega_{\text{C}}\frac{1}{N_{\text{C}}}\sum_{i=1}^{N_{\text{C}}}|c_{i}-\tilde{c}_{i}|,$ (11)
-where $\omega_{\text{L}}$, $\omega_{\text{sur}}$ and $\omega_{\text{C}}$ are weighting parameters to balance the contributions from the three losses.
-$s_{i}$ is the predicted signed distance value at the observed point along the sampled ray.
+$\mathcal{L}_{\text{rend}}=\frac{1}{N_{\text{L}}}\sum_{i=1}^{N_{\text{L}}}(|r_{i}-\tilde{r}_{i}|+\omega_{\text{sur}}|s_{i}|)+\frac{\omega_{\text{C}}}{3\cdot N_{\text{C}}}\sum_{i=1}^{N_{\text{C}}}\sum_{j=1}^{3}|c_{i}^{j}-\tilde{c}_{i}^{j}|,$ (10)
+where $\omega_{\text{sur}}$ and $\omega_{\text{C}}$ are weighting parameters for the losses.
+$r_{i}$ is the observed range along the $i^{th}$ sampled ray and $s_{i}$ is the predicted signed distance value at the observed points.
+$c_{i}^{j}$ is the value of $j^{th}$ channel in the image pixel, where $j=\{1,2,3\}$ corresponds to RGB channels.
 Figure 3: Inspiration of Curvature Sampling.
-From the two zoom-in areas, it can be found that areas with low curvature tend to be less informative, like road plane, while those with high curvature might provide more information for pre-training (objects like cars).
+It can be found that areas with low curvature tend to be less informative, like road plane, while those with high curvature provide more information (objects like cars).
 3.3 Curvature Sampling
 In order to make joint unsupervised representation learning feasible, we have to make $N_{\text{L}}\ll N_{\text{P}}$ and $N_{\text{C}}\ll H\cdot W\cdot N_{\text{cam}}$.
-Intuitively, uniform sampling can be used.
-But due to the relatively small sample number compared to the raw inputs ($\sim\frac{1}{100}$), uniform sampling brings little improvement against separate pre-training, which is contradictory to our motivation.
+Intuitively, uniform sampling with depth can be used, same as “Memory-friendly Ray Sampling” in [58].
+But due to the relatively small sample number compared to the raw inputs ($\sim\frac{1}{100}$), this sampling method brings little improvement against separate pre-training, which is contradictory to our motivation.
 Hence, we need to sample more informative part of the scene for $\mathcal{L}_{\text{rend}}$.
@@
 $\mathbf{n}=\frac{\delta f^{\text{SDF}}([\mathbf{p},\mathbf{f}])}{\delta\mathbf{p}},$ (11)
-where $\mathbf{n}\in\mathbb{R}^{3}$ is the predicted normal. Then we normalized $\mathbf{n}$ to get the direction of the normal
-$\tilde{\mathbf{n}}=\frac{\mathbf{n}}{||\mathbf{n}||_{2}}.$ (13)
+where $\mathbf{n}\in\mathbb{R}^{3}$ is the predicted normal.
+Then we normalized $\mathbf{n}$ to get the direction of the normal $\tilde{\mathbf{n}}=\frac{\mathbf{n}}{||\mathbf{n}||_{2}}$.
 Here $||\cdot||_{2}$ is the L-2 norm.
-Next we estimate the mean curvature $\mathbf{c}\in\mathbb{R}^{3}$ at $\mathbf{p}$ by applying the vector differential operator on $\tilde{\mathbf{n}}$ with respect to $\mathbf{p}$
-$\mathbf{c}=\nabla_{\mathbf{p}}\tilde{\mathbf{n}}.$ (14)
-For each point $\mathbf{p}_{n}$ in LiDAR point cloud, we compute the average of $\mathbf{c}_{n}$ as the sampling weights $\omega_{n}$ and sample $N_{\text{L}}$ LiDAR point cloud with a Multinomial Sampler implemented in PyTorch [35] for differentiable rendering.
-For pixels on image plane, we project the LiDAR point cloud back to image planes with $\mathcal{T}$, assign $\omega_{n}$ of each point to the projected pixel and apply a gaussian blur kernel of size $K_{\text{gaus}}$ to densify the weights, with which we sample $N_{\text{C}}$ pixel for $\mathcal{L}_{\text{rend}}$.
-As the estimation of curvature here is noisy especially during the first few epochs, we set up a warm-up stage with $N_{\text{warmup}}$ epochs with uniform sampling and after $N_{\text{warmup}}$ epochs, Curvature Sampling is utilized.
-3.4 Swapping Prototype Assignment Prediction
-CLAP aims to utilize the interaction between camera and LiDAR encoders for joint unsupervised 3D representation learning.
-Meanwhile, we hope to learn a common feature space standing for segments of different objects in an unsupervised manner.
-Inspired by SwAV [3], we propose to use learnable prototypes to represent parts of the scenes.
-We first randomly initialize $N_{\text{K}}$ learnable prototypes $\mathbf{K}\in\mathbb{R}^{N_{\text{K}}\times d_{\text{K}}}$, each of which is a $d_{\text{K}}$ vectors.
-Then we pre-train the backbones to understand the interaction between different modalities with a swapping prototype assignment prediction loss [3].
-Furthermore, we introduce an Expectation-Maximization scheme to optimize the learnable prototypes and a gram matrix minimization loss to avoid collapse [3] of the prototype learning.
-Swapping Prototype Assignment Prediction.
-We project the LiDAR embeddings $\hat{\mathbf{P}}$ and camera embeddings $\hat{\mathbf{I}}$ separately with two projection heads $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ to the same dimension as $\mathbf{K}$
+Next we apply the differential operator on $\tilde{\mathbf{n}}$ with respect to $\mathbf{p}$ and get $\mathbf{c}\in\mathbb{R}^{3}$.
+$\mathbf{c}=\frac{\delta\tilde{\mathbf{n}}}{\delta\mathbf{p}}.$ (12)
+Then we estimate the geodesic curvature [48] of each point $\mathbf{p}_{n}$ in LiDAR point cloud by computing the norm of $\mathbf{c}_{n}$ and use it as the sampling weights $\omega_{n}$, that is $\omega_{n}=||\mathbf{c}_{n}||_{2}$.
+With $\omega_{n}$, we sample $N_{\text{L}}$ points with a Multinomial Sampler implemented in PyTorch [37] for differentiable rendering.
+For pixels on image plane, we project the LiDAR point cloud back to image planes with $\mathcal{T}$, assign $\omega_{n}$ of each point to the projected pixel and apply a gaussian blur kernel of size $K_{\text{gaus}}$ to densify the weights, with which we sample $N_{\text{C}}$ pixels for $\mathcal{L}_{\text{rend}}$.
+As the estimation of curvature here is noisy especially in the first few epochs, we set up a warm-up stage with $N_{\text{warmup}}$ epochs with uniform sampling and after $N_{\text{warmup}}$ epochs, Curvature Sampling is utilized.
+3.4 Prototype Learning
+Curvature Sampling allows for joint pre-training, prompting us to explore how camera and LiDAR data can be used to understand “objectness” or “object parts” in an unsupervised way.
+We use learnable prototypes to represent segments of 3D scenes and establish a shared feature space that connects the two modalities.
+Firstly, we randomly initialize $N_{\text{K}}$ learnable prototypes $\mathbf{K}\in\mathbb{R}^{N_{\text{K}}\times d_{\text{K}}}$, each of which is a $d_{\text{K}}$ vectors.
+Then an Expectation-Maximization training approach is proposed to maximize the similarity between these prototypes and 3D embeddings from the two modalities.
+To delve deeper into the interplay between image semantics and LiDAR geometry, we use the swapping prototype prediction loss.
+Finally, to avoid the prototype collapsing into one same vector, we introduce a Gram Matrix Regularization Loss.
+Expectation-Maximization.
+In order to guide the learnable prototypes to represent parts of the environment, we propose an Expectation-Maximization [35] training scheme to optimize the prototypes.
+We first project the LiDAR embeddings $\hat{\mathbf{P}}$ and camera embeddings $\hat{\mathbf{I}}$ separately with two projection heads $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ to the same dimension as $\mathbf{K}$
 $\dot{\mathbf{P}}=f^{\text{proj}}_{\text{P}}(\hat{\mathbf{P}})\ \ ,\ \ \dot{\mathbf{I}}=f^{\text{proj}}_{\text{I}}(\hat{\mathbf{I}}),$ (13)
-where $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ are parameterized by Multi-Layer Perceptron and $\dot{\mathbf{P}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{K}}}$, $\dot{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{K}}}$.
-We denote $N_{\text{3D}}=\hat{D}\times\hat{H}\times\hat{W}$ and then normalize and reshape the projected embeddings into $\dot{\mathbf{P}}\in\mathbb{R}^{N_{\text{3D}}\times\hat{d}_{\text{K}}}$ and $\dot{\mathbf{I}}\in\mathbb{R}^{N_{\text{3D}}\times\hat{d}_{\text{K}}}$.
+where $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ are parameterized by Multi-Layer Perceptron and $\dot{\mathbf{P}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times d_{\text{K}}}$, $\dot{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times d_{\text{K}}}$.
+We denote $N_{\text{3D}}=\hat{D}\times\hat{H}\times\hat{W}$ and then normalize and reshape the projected embeddings into $\dot{\mathbf{P}}\in\mathbb{R}^{N_{\text{3D}}\times d_{\text{K}}}$ and $\dot{\mathbf{I}}\in\mathbb{R}^{N_{\text{3D}}\times d_{\text{K}}}$.
 After that, similarity scores $\mathbf{S}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$ between 3D embeddings and prototypes are computed separately for LiDAR and camera branches
 $\mathbf{S}_{\text{P}}=\dot{\mathbf{P}}\cdot\mathbf{K}^{\intercal}\ \ ,\ \ \mathbf{S}_{\text{I}}=\dot{\mathbf{I}}\cdot\mathbf{K}^{\intercal}.$ (14)
-Similar to SwAV [3], we detach $\mathbf{S}_{\text{P/I}}$ and apply sinkhorn algorithm [9] to make them approach the double stochastic matrix in $N_{\text{sink}}$ iterations.
-We denote the updated matrix as codes $\mathbf{Q}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$.
-The swapping prototype assignment loss is formulated with a temperature parameter $\tau$ [3],
-$\begin{split}\mathcal{L}_{\text{SwAV}}&=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\mathbf{Q}_{\text{I}}^{n,m}\log\frac{exp(\mathbf{S}_{\text{P}}^{n,m})/\tau}{\sum_{k=1}^{N_{\text{K}}}exp(\mathbf{S}_{\text{P}}^{n,k})/\tau}\\ &+\mathbf{Q}_{\text{P}}^{n,m}\log\frac{exp(\mathbf{S}_{\text{I}}^{n,m})/\tau}{\sum_{k=1}^{N_{\text{K}}}exp(\mathbf{S}_{\text{I}}^{n,k})/\tau}\}.\end{split}$ (17)
-Expectation-Maximization.
-In order to guide the learnable prototypes to represent parts of the environment, we introduce the Expectation-Maximization algorithm [33] to further optimize the prototypes.
 In the Expectation step, we compute the probability $\hat{\mathbf{S}}_{\text{P/I}}$ that each prototype is assigned to each embeddings by applying a softmax operation on $\mathbf{S}_{\text{P/I}}$.
@@
 $\mathcal{L}_{\text{EM}}=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\hat{\mathbf{S}}_{\text{P}}^{n,m}log\hat{\mathbf{S}}_{\text{P}}^{n,m}+\hat{\mathbf{S}}_{\text{I}}^{n,m}log\hat{\mathbf{S}}_{\text{I}}^{n,m}\}$ (15)
+Swapping Prototype Prediction.
+To further explore interaction between modalities, we detach $\mathbf{S}_{\text{P/I}}$ and apply sinkhorn algorithm [9] to make them approximate to double stochastic matrix in $N_{\text{sink}}$ iterations.
+We denote the updated matrix as codes $\mathbf{Q}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$.
+The swapping prototype prediction loss is formulated with a temperature parameter $\tau$, inspired by [4],
+$\begin{split}\mathcal{L}_{\text{SwAV}}&=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\mathbf{Q}_{\text{I}}^{n,m}\log\frac{exp(\mathbf{S}_{\text{P}}^{n,m})/\tau}{\sum_{k=1}^{N_{\text{K}}}exp(\mathbf{S}_{\text{P}}^{n,k})/\tau}\\ &+\mathbf{Q}_{\text{P}}^{n,m}\log\frac{exp(\mathbf{S}_{\text{I}}^{n,m})/\tau}{\sum_{k=1}^{N_{\text{K}}}exp(\mathbf{S}_{\text{I}}^{n,k})/\tau}\}.\end{split}$ (16)
 Gram Matrix Minimization.
@@
 Rand. 48.69 55.28 78.52 46.64 16.18 50.44 22.11 57.00 46.87 30.56 76.16 62.40
-UniPAD 49.81 +0.72 55.29 +0.01 80.81 42.81 17.08 48.98 25.85 61.72 50.19 27.53 78.53 64.57
+ALSO 45.34 -3.35 49.19 -6.09 77.61 41.98 14.61 36.09 12.55 58.10 42.89 30.56 74.09 64.98
+OCC-MAE 47.39 -1.30 54.97 -0.31 77.43 46.25 15.26 50.06 19.35 55.35 43.24 30.92 74.69 61.34
+SLidR 47.23 -1.46 52.77 -2.51 77.12 45.04 16.19 50.50 22.17 57.74 41.47 30.22 71.11 60.71
+PPKT 49.58 +0.89 55.85 +0.57 79.24 47.26 17.26 51.37 21.14 59.55 44.82 31.43 78.03 65.73
+UniPAD 49.81 +1.12 55.29 +0.01 80.81 42.81 17.08 48.98 25.85 61.72 50.19 27.53 78.53 64.57
 CLAP 51.17 +2.48 57.04 +1.76 79.56 48.43 18.84 56.34 23.98 60.60 48.87 34.11 78.08 62.87
 Table 1: Results for fine-tuning on 5% of training set in NuScenes [3].
-“Init.” means the way to initialize models, including the baseline UniPAD [53].
+“Init.” means the way to initialize models.
 To avoid the case the performance difference is brought by accelerating convergence, we first increase the training iteration of training-from-scratch model (BEVFusion [30]) until we observe convergence, which leads to the “Rand.” row.
-Then we pre-train the backbones with UniPAD [53] and the proposed CLAP, initialize BEVFusion with the pre-trained weights and then fine-tune the model with the same number iterations as “Rand.”.
+Then we pre-train the backbones with ALSO [2], OCC-MAE [34], SLidR [38], PPKT [28], UniPAD [58] and CLAP.
+We initialize BEVFusion with the pre-trained weights and fine-tune the model with the same number iterations as “Rand.”.
 We provide mAP and NDS as an evaluation of the overall performance of different models and highlight the best mAP and NDS with bold font.
 We also indicate the performance improvement by green color.
-For detailed categories, we provide AP for them and “C.V.”, “Mot.”, “Bic.”, “Ped.” and “T.C.” are abbreviations for Construction Vehicle, Motorcycle, Bicycle, Pedestrian and Traffic Cone.
+“C.V.”, “Mot.”, “Bic.”, “Ped.” and “T.C.” are abbreviations for Construction Vehicle, Motorcycle, Bicycle, Pedestrian and Traffic Cone.
 All the results here are in %.
@@
 Unsupervised 3D representation learning for fusion perception aims to pre-train both LiDAR and camera encoders and initialize downstream models with the pre-trained weights to gain performance improvement in downstream tasks.
-In this section, we design extensive experiments on the popular autonomous driving dataset NuScenes [2] to demonstrate the effectiveness of CLAP.
+In this section, we design extensive experiments on the popular autonomous driving dataset NuScenes [3] and Waymo [45] to demonstrate the effectiveness of CLAP.
 To begin with, we describe experiment setups in Section 4.1.
@@
 Datasets.
-We use the popular autonomous driving dataset NuScenes [2] to evaluate the performance of CLAP.
+We use the popular autonomous driving dataset NuScenes [3] and Waymo [45] to evaluate the performance of CLAP.
 NuScenes [3] uses one roof LiDAR and six surrounding cameras to collect data.
@@
 Following convention practice from [3, 46], we divide the whole dataset into training set with 850 scenes and validation set with 150 scenes.
-Pre-trainings are conducted on the whole training set and downstream task is 3D object detection under few-shot setting, where we randomly sample $5\%$ labeled data in the training set to train the randomly initialized model and fine-tune the pre-trained ones.
-The overall mAP (mean accurate precisions) and NDS (NuScenes Detection Score) are reported, along with AP (accurate precisions) for different categories.
+Waymo [45] uses one top 64-beam LiDAR, 4 corner LiDARs and 5 surrounding cameras to collect point clouds and camera images.
+Following the same practice in [45, 46], the collected 1000 scenes are split into training set (798 scenes) and validation set (202 scenes).
+All pre-trainings are conducted on the training set without labels and we conduct downstream 3D object detection training in few-shot setting via uniform sampling of the training data (NuScenes [3] $5\%$ and Waymo [45] $1\%$).
 Downstream 3D Object Detectors.
-We conduct experiments on one of the SOTA fusion 3D object detector called BEVFusion [27].
-BEVFusion is implemented in the popular code repository for autonomous driving perception called OpenPCDet [42].
-We use average precisions of various categories (APs), mean average precision (mAP) and NuScenes Detection Score (NDS) [2] as the evaluation metrics.
-We follow a similar setting in [48] to gradually increase training iterations of the model without pre-trained weights until convergence is observed.
+For NuScenes [3], we select the SOTA 3D object detector called BEVFusion [30] and for Waymo [45], we use CenterPoint [59].
+Both BEVFusion and CenterPoint are implemented in the popular code repository for autonomous driving perception called OpenPCDet [46].
+For evaluation metrics, we use average precisions of various categories (APs), mean average precision (mAP) and NuScenes Detection Score (NDS) [3] for NuScenes [3] and mAP (mean accurate precisions) and mAPH (mean accurate precisions with headings) at different difficulty levels (Level-1 and 2) for Waymo [45].
+We follow a similar setting in [14, 53] to gradually increase training iterations of the from-scratch model until convergence is observed.
 Here convergence means further increasing training iterations will not improve the performance.
 Then the number of training iterations is fixed for fine-tuning pre-trained models.
-This setting avoid the case that pre-training only accelerate convergence and make sure that pre-training indeed improve the performance of downstream models, that is improving the sample efficiency of the downstream task.
+This setting avoids the case that pre-training only accelerates convergence and makes sure that pre-training indeed improve the performance of downstream models, that is improving the sample efficiency of the downstream task.
 Baseline Pre-training Method for Fusion Perception.
-We incorporate UniPAD [53] as the baseline 3D unsupervised pre-training method.
-We utilize the official implementation from [53] to pre-train the backbones of BEVFusion [27].
-As pre-trained image backbones are broadly used for image feature extraction, the implementation of [53] also use a pre-trained image backbone from [8] to initiate their pre-training and the training set for this backbone does not involve any data from NuScenes [2].
-We adopt this practice and also for training-from-scratch model, we initialize the camera encoder with the same pre-trained weights from [8].
+We incorporate three kinds of pre-training baseline methods: 1) an occupancy estimation method called ALSO [2], 2) occupancy masked autoencoder called Occupancy-MAE [34], 3) multi-modality methods including SLidR [38], PPKT [28] and UniPAD [58].
+We use the official implementations to pre-train the backbones with same setting as CLAP.
 Implementation Details of CLAP.
-The feature channels for LiDAR embeddings $\hat{\mathbf{P}}$, image embeddings $\hat{\mathbf{I}}$, fusion embeddings $\hat{\mathbf{F}}$ and swapping prototype assignment prediction are respectively set to $\hat{d_{\text{P}}}=256$, $\hat{d_{\text{I}}}=80$, $\hat{d_{\text{F}}}=512$ and $\hat{d}_{\text{K}}=128$.
-Sampling number for LiDAR point cloud and camera pixel are $N_{\text{L}}=8192$ and $N_{\text{C}}=1024\times N_{\text{cam}}$, where $N_{\text{cam}}=6$ for NuScenes dataset [2].
+The feature channels for embeddings $\hat{\mathbf{P}}$, $\hat{\mathbf{I}}$, $\hat{\mathbf{F}}$ and prototypes $\mathbf{K}$ are respectively set to $\hat{d_{\text{P}}}=256$, $\hat{d_{\text{I}}}=80$, $\hat{d_{\text{F}}}=512$ and $\hat{d}_{\text{K}}=128$.
+Sampling number for point cloud and pixel are $N_{\text{L}}=8192$ and $N_{\text{C}}=1024\times N_{\text{cam}}$.
 The number of sample points along each ray is $N_{\text{ray}}=96$.
 Warm-up epochs for Curvature Sampling is $N_{\text{warmup}}=4$.
-For swapping prototype assignment prediction, we set the number of learnable prototypes and sinkhorn update iterations to $N_{\text{K}}=512$ and $N_{\text{sink}}=3$.
+We set the number of learnable prototypes and sinkhorn update iterations to $N_{\text{K}}=512$ and $N_{\text{sink}}=3$.
 The temperature for swapping prediction loss is $\tau=1.0$.
-The loss weighting parameters are implemented as $\omega_{\text{proto}}=1.0$, $\omega_{\text{L}}=2.0$, $\omega_{\text{sur}}=0.1$, $\omega_{\text{C}}=0.1$, $\omega_{\text{SwAV}}=1.0$, $\omega_{\text{EM}}=0.1$ and $\omega_{\text{GMM}}=0.1$.
-We use a learning rate of 0.00005 with a cosine learning schedule for pre-training and use mask augmentation for CLAP with a masking rate of 0.9.
+The loss weighting parameters are implemented as $\omega_{\text{r}}=2.0$, $\omega_{\text{proto}}=1.0$, $\omega_{\text{sur}}=0.05$, $\omega_{\text{C}}=0.05$, $\omega_{\text{SwAV}}=1.0$, $\omega_{\text{EM}}=0.1$ and $\omega_{\text{GMM}}=0.1$.
+We use torch.auto_grad() [37] to implement the derivatives in the curvature estimation.
+For more details about pre-training and fine-tuning, please refer to Appendix 1.
 4.2 Main Results
-Both CLAP and baseline methods are pre-trained on the whole training set of NuScenes [2] without labels.
-To compare the effectiveness of CLAP and UniPAD [53], we randomly sample 5% of the training labels and fine-tune BEVFusion [27] initialized by the pre-trained weights.
-Furthermore, as the size of pre-training dataset is limited, we randomly sample 2.5%, 1% and 0.5% of the training labels and fine-tune BEVFusion with CLAP pre-trained weights to see whether potential scaling property exists in pre-training with CLAP.
-Comparison to Previous SOTA Method.
-As shown in Table 1, CLAP achieves 2.48% mAP improvement over randomly initialization at convergence, which is $300\%$ more improvement for mAP than SOTA unsupervised 3D representation method UniPAD [53].
-For NDS metric, UniPAD [53] only achieves comparable performance while CLAP surpasses the train-from-scratch model by 1.76% .
+NuScenes Results.
+As shown in Table 1, CLAP achieves 2.48% mAP improvement over randomly initialization at convergence, which is $100\%$ more improvement for mAP than SOTA unsupervised 3D representation method UniPAD [58] and the best among all initialization methods.
+For NDS metric, UniPAD [58] only achieves comparable performance and PPKT [28] makes a gain of 0.57% while CLAP surpasses the train-from-scratch model by 1.76% .
 When it turns to different categories, CLAP generally benefit the performance of all the categories and for Construction Vehicle, Bus, Barrier, Motorcycle and Bicycle, the improvement over random initialization are more than 2% AP.
+Init. Level-1 Level-2 $\bar{\Delta}$
+mAP mAPH mAP mAPH
+Rand. 61.60 58.58 55.62 52.87 0
+ALSO 62.09 59.03 56.12 53.32 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.47}}$
+OCC-MAE 62.33 59.32 56.36 53.63 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.74}}$
+SLidR 62.10 59.09 56.10 53.36 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.49}}$
+PPKT 62.32 59.22 56.37 53.55 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.69}}$
+UniPAD 61.57 58.64 55.64 52.93 $\color[rgb]{0.1328125,0.546875,0.1328125}{{+0.02}}$
+CLAP 62.87 59.88 56.88 54.16 $\color[rgb]{0.1328125,0.546875,0.1328125}{\mathbf{+1.28}}$
+Table 2: Results for fine-tuning on $1\%$ of training set in Waymo [45].
+We first increase training iterations of from-scratch model and observe convergence (Rand.).
+Then we fix the training schedule to fine-tune the pre-training models.
+We provide an average difference $\bar{\Delta}$ to show the performance gain.
+All results are in %.
+Waymo Results.
+In Table 2, we provide the average performance difference of mAP and mAPH over different difficulty levels.
+It can be found that CLAP achieves the best performance at convergence.
+Meanwhile, the performance gain brought by CLAP is approximately two times as the best (OCC-MAE [34]) of previous pre-training methods.
+This demonstrates the effectiveness and generalization ability of CLAP.
 Potential Scaling Property.
-As we are not able to scale up the pre-training dataset at current stage, we explore potential scaling property by gradually decreasing the sample numbers (2.5%, 1% and 0.5%) for fine-tuning, which increases the ratio between pre-training data and fine-tuning data.
-The results are shown in Table 2 together with the results of fine-tuning on 5% of training set in NuScenes.
+As we are not able to scale up the pre-training dataset at current stage, we explore potential scaling property by gradually decreasing the sample numbers (2.5%, 1% and 0.5%) for fine-tuning on NuScenes, which increases the ratio between pre-training data and fine-tuning data.
+The results are shown in Table 3.
 It can be found that as the ratio between pre-training data and fine-tuning data gets larger, the performance improvement by CLAP increases and CLAP provides a gain up to 7.22% mAP and 4.71% NDS with 0.5% fine-tuning data.
@@
 Table 3: Results on potential scaling property of CLAP .
-“Init.” means the method for initialization.
-“Fine-tune Data” stands for the sample number of training set in NuScenes [2] used for fine-tuning/training-from-scratch.
-We gradually increase the number of training iterations for training-from-scratch model (BEVFusion [27]) with different numbers of training data until we observe convergence, leading to the results of “Rand.”.
-Then we initialize BEVFusion with weights pre-trained by CLAP and then fine-tune the model with the same numbers of iterations as “Rand.”.
-We provide mAP and NDS as an evaluation of the overall performance.
-All the results here are in %.
+“Fine-tune Data” stands for the sample number of training set in NuScenes [3].
+Results for random initialization are at convergence and fine-tuning with CLAP uses the same numbers of iterations.
+Results are mAP and NDS in %.
 Figure 4: Visualization of curvature estimation.
@@
 4.3 Visualizations
-In this section, we provide visualization to show the effectiveness of the proposed Curvature Sampling.
-In order to evaluate whether the estimated curvature is able to help sampling more informative part of the environment, we use the pre-trained model to estimate the curvature of LiDAR point clouds and use heatmap color to indicate the weight computed in Section 3.3.
+We use the pre-trained model to estimate the curvature of LiDAR point clouds and use heatmap color to indicate the weight computed in Section 3.3.
 The visualization results are shown in Figure 4.
 We use orange boxes to highlight those regions with relatively correct estimation and green ones for those with noisy estimation.
-Furthermore, for better understanding about the curvature estimation, we zoom in for some correctly estimated regions.
-It can be found that though some noise exists, CLAP is able to predict high weights for those highly informative region for sampling and meanwhile assign lower weights to most of the background, which makes joint pre-training feasible.
+Furthermore, we zoom in some orange boxes.
+It can be found that though some noise exists (because pre-training is conducted in unsupervised manner), CLAP is able to predict high weights for those highly informative region for sampling and meanwhile assign lower weights to most of the background, which makes joint pre-training feasible.
 Joint Pre-train Cur. Sam. Proto. Learning mAP
@@
 We conduct ablation study to evaluate the effectiveness of different components.
-As shown in Table 3, it can be found that using uniform sampling (second line) to make joint pre-training feasible does not bring improvement over separate pre-training (first line) [53].
-Then, Curvature Sampling (third line) improves the performance over uniform sampling and separate pre-training by assigning larger weights to more informative parts of the 3D scene during sampling.
-Finally, the Prototype Learning scheme (fourth line) introduces interaction of LiDAR and camera encoders into pre-training and achieve the best performance.
+Results are in Table 4.
+The first line is seperate pre-training with UniPAD [58].
+The second one is jointly pre-training using UniPAD [58] and uniform sampling guided by depth (“Memory-friendly Ray Sampling” in [58]) to address the GPU memory limitation.
+It can be found that using simple sampling method does not bring improvement over separate pre-training.
+Then we add Curvature Sampling in third line and found that Curvature Sampling enables more effective sampling and improves the performance over uniform sampling guided by depth and separate pre-training.
+Finally, the Prototype Learning scheme (fourth line) learns a common feature space for segments of 3D scenes and introduces interaction of LiDAR and camera encoders, which achieves the best performance.
 5 Conclusion
-In this paper, we propose CLAP to jointly pre-train for fusion perception in an unsupervised manner via differentiable rendering.
-To make joint pre-training feasible, CLAP utilizes the novel Curvature Sampling to sample more informative parts of the 3D scene for pre-training.
-Further, a swapping prototype prediction scheme is introduced in order to make full use of the interaction between LiDAR point clouds and camera images.
-To make the learnable prototypes better represent parts/semantics of objects, CLAP brings the idea of Expectation-Maximization for prototypes optimization, alongside with a gram matrix regularization term to avoid collapse.
+In this paper, we propose CLAP for unsupervised fusion joint pre-training via differentiable rendering.
+CLAP uses Curvature Sampling to sample more informative parts and learnable prototypes to represent parts of 3D scenes, optimized with an Expectation-Maximization approach.
+To further make use of the interplay between LiDAR geometry and image semantics, a swapping prediction loss is used.
+Finally, a Gram Matrix Regularization loss is proposed to avoid collapse.
 Experiment results demonstrate that CLAP is superior in unsupervised 3D representation learning and has the potential to scale up.
-In the future, when a larger pre-training dataset is available, we expect to scale up the pre-training with CLAP for more performance improvement.
-Supplementary Material
-Init. F.T. mAP Vehicle Pedestrian Cyclist
-0-30m 30-50m 50m- 0-30m 30-50m 50m- 0-30m 30-50m 50m-
-Rand* 5% 20.48 58.03 25.22 12.98 11.62 9.75 6.97 21.55 6.83 3.11
-CLAP* 22.86 +2.38 58.37 26.38 14.07 12.60 9.50 7.88 30.08 10.50 5.39
-Rand 46.07 76.71 51.15 31.84 37.53 20.12 9.84 62.00 42.61 24.18
-CLAP 46.88 +0.81 76.98 51.64 31.31 38.79 20.60 9.74 63.75 43.21 26.83
-Rand* 100% 64.00 86.21 70.20 58.20 57.80 41.18 23.55 75.95 61.45 45.80
-CLAP* 64.74 +0.74 88.14 72.59 59.13 57.37 42.24 24.22 77.11 61.91 45.63
-Rand 65.03 88.18 74.23 61.75 57.32 38.90 21.96 78.07 64.32 48.16
-CLAP 65.56 +0.53 87.97 72.77 62.11 58.33 40.11 21.29 78.63 64.70 47.27
-Table 4: Results for transferring experiments on Once [29] dataset.
-CenterPoint [54] is used as the downstream detector.
-“Init.” indicates the initialization methodshorthands.
-“F.T.” indicates the number of training samples in fine-tuning stage.
-Overall mAP and APs for different categories within different ranges are shown in this table.
-“Rand*” means training the randomly initialized model with the original training iterations in OpenPCDet [42].
-“Rand” indicates that we increase the number of training iterations for randomly initialized model until convergence is observed.
-“CLAP*” indicates that we pre-train the backbones with CLAP on NuScenes [2] and then fine-tune on Once with the original iterations in [42].
-“CLAP” uses the same fine-tuning iterations as “Rand”.
-We use green color to highlight the performance improvement brought by CLAP.
-All the results are in %.
-A Transferring to other datasets
-We conduct further experiments to evaluate the transferring ability of CLAP.
-Specifically, we select LiDAR-based 3D object detector CenterPoint [54] on Once [29] for the downstream task.
-A 40-beam LiDAR is utilized in Once [29] to collect 15k labeled training data.
-We randomly sample $5\%$ and also use all of the labeled training set to train the from-scratch model until convergence is observed.
-Then we use pre-trained weights by CLAP on NuScenes [2] to initialize the same model and fine-tune it with the same training iterations as the randomly initialized model.
-Results are shown in Table 4.
-It can be found that pre-training by CLAP also benefits LiDAR-based 3D object detection, even in a cross-dataset setting.
-And if we look at the performance of “Rand*” and “CLAP*”, CLAP also accelerates the convergence in downstream task.
-B Visualization of Prototype Learning
-We use the model pre-trained by CLAP to infer the 3D features and assign prototypes to different LiDAR points in the 3D space.
-Then we use different colors to indicate different prototypes and visualize them in Figure 5.
-If can be found that the background road plane inside the same frame is generally assigned to the same prototype.
-And foreground vehicles are assigned to another prototype.
-This demonstrate that the proposed prototype learning scheme actually learns some level of semantics understanding in an unsupervised manner.
-However, as our pre-training does not incorporate any label, it can also be found that the prototype assignment has some noise, for example some of the road plane points are assigned to other prototypes.
-Figure 5: Visualization of the prototype learning results. Different color indicates different prototype assignments.

### v3

@@
-CLAP: Unsupervised 3D Representation Learning for Fusion 3D Perception via Curvature Sampling and Prototype Learning
+arXiv:2412.03059v3[cs.CV] 27 Feb 2026
 Abstract
@@
 Experiments on NuScenes and Waymo datasets show that CLAP achieves up to $100\%$ more performance gain as compared to previous SOTA pre-training methods.
-Codes and models will be released.
+Codes and models will be available here.
 1 Introduction
@@
 The autonomous system, on which these applications are deployed, are typically equipped with multiple sensors including visual ones like cameras that produce RGB images, and range sensors like LiDAR (Light-Detection-And-Ranging) that generate point clouds.
-Fusion of these two modalities [30, 26, 22, 25, 21, 32, 8] have generally improved over use of a single modality, e.g., camera [11, 7, 43, 50, 60, 20] or LiDAR [39, 55, 42, 59, 12, 1] separately, in 3D perception performance.
-Figure 1: Unlike previous SOTA unsupervised 3D representation learning method UniPAD [58] that separately pre-train LiDAR and camera encoders with differential rendering (a), our proposed method CLAP conducts joint pre-training for fusion perception.
+Fusion of these two modalities (Liu et al., 2023; Liang et al., 2022; Li et al., 2022b; Liang et al., 2019; Li et al., 2022a; Meyer et al., 2019; Chen et al., 2017) have generally improved over use of a single modality, e.g., camera (Ding et al., 2020; Chen et al., 2016; Simonelli et al., 2019; Wang et al., 2021b; Zhang et al., 2021; Ku et al., 2019) or LiDAR (Shi et al., 2019; Yan et al., 2018; Shi et al., 2021; Yin et al., 2021; Fan et al., 2021; Bai et al., 2022) separately, in 3D perception performance.
+Figure 1: Unlike previous SOTA unsupervised 3D representation learning method UniPAD (Yang et al., 2024) that separately pre-train LiDAR and camera encoders with differential rendering (a), our proposed method CLAP conducts joint pre-training for fusion perception.
 However, training multimodal 3D perception models is expensive as labeling in 3D space is notoriously time-and-energy-consuming.
 Unsupervised 3D representation learning, which pre-trains backbones without any label and fine-tuned the pre-trained weight for downstream performance improvement, has shown the potential to alleviate the labeling burden in 3D perception.
-Amongst the many unsupervised 3D pre-training methods [24, 19, 6, 57, 53, 18, 62, 58], use of mask auto-encoding (reconstruction) and differntiable rendering has emerged as the most performant.
-However, encoding and processing high-dimensional multi-modal data (images and point clouds) is computationally expensive: If one were to pre-train using all points and pixels within the point cloud and image, respectively, even the most advanced GPU to date (H100) is only able to hold a batch size of 1.
-Therefore, existing methods, e.g., UniPad [58], has conventionally pre-trained each modality separately.
-While the two modalities are complementary (semantics from image, and structure from point clouds), current methods are unable to exploit their mutual benefits during pre-training.
-To address this pain point, we propose a joint unsupervised pre-training method of image and point cloud modalities based on differentiable rendering.
+Amongst the many unsupervised 3D pre-training methods (Liang et al., 2021; Huang et al., 2021; Chen et al., 2022; Yang et al., 2023; et al, 2023; Huang et al., 2023; Zhu et al., 2023; Yang et al., 2024), use of mask auto-encoding (reconstruction) and differntiable rendering has emerged as the most performant.
+However, encoding and processing high-dimensional multi-modal data (images and point clouds) is computationally expensive: If one were to pre-train using all points and pixels within the point cloud and image, even the most advanced GPU to date is only able to hold a batch size of 1.
+Therefore, existing methods, e.g., UniPad (Yang et al., 2024), has conventionally pre-trained each modality separately.
+The limitation of separate pre-training is that each encoder is restricted to its own modality.
+Recovering 3D information from images is an ill-posed problem; point clouds provide geometry cues but lack higher-level semantics.
+To address this pain point, we propose a joint unsupervised pre-training method of image and point cloud modalities based on differentiable rendering to exploit their complementarity.
 Our method, CLAP, short for Curvature sampLing and leArnable Prototype, addresses the primary computational challenge by a curvature sampling strategy, which comes from the observation that there exists redundancy in information when sampling multiple points on the same flat surface.
 This is enabled by estimating the curvature of each point in the 3D space by taking second order derivative of the SDF (signed distance field) function.
-This allows us to identify points at which large change in the surface occurs.
-Through calibration, we can obtain the corresponding pixel within the image, around which we sample assuming the coordinates are normally distributed.
-This sampling strategy captures variations in the point cloud and in turn provide more informative points than the conventional sampling strategy in [58].
+This sampling strategy captures variations in the point cloud and in turn provide more informative points than the conventional strategy in (Yang et al., 2024).
 As our curvature sampling strategy has reduced the computational burden to allow both modalities to be simultaneously processed, we propose to model the interplay between image and point cloud modalities through a set of learnable prototypes, which represents parts of the 3D scene and enables a common feature space to bridge the two modalities.
@@
 Furthermore, we propose to use the swapping prediction loss to explore the interaction between the two modalities.
-Last but not least, we utilize a Gram Matrix Regularization term to minimize similarity across prototypes, which addresses collapse of prototype when trained naively.
+Last but not least, we utilize a Gram Matrix Regularization term to minimize similarity across prototypes, avoiding collapse of prototype when trained naively.
 Our contributions are 4-folds: (1) We propose a curvature sampling strategy to identify informative points (and pixels) for sampling, which enables the first joint differentiable-rendering-based pre-training method for fusion perception.
@@
 (3) We further propose to use a swapping prediction loss for modality interplay and a Gram Matrix Regularization loss that avoids collapse of prototype learning.
-(4) Through extensive experiments on the popular autonomous driving datasets NuScenes [3] and Waymo [45], we demonstrate the effectiveness of CLAP .
+(4) Through extensive experiments on the popular autonomous driving datasets NuScenes (Caesar et al., 2020) and Waymo (Sun et al., 2020), we demonstrate the effectiveness of CLAP .
 For example, CLAP achieves up to $100\%$ more improvement than previous SOTA 3D pre-training methods and shows potential scaling property.
@@
 For LiDAR-based 3D object detection, there are three main streams with different embedding schemes for point clouds inputs.
-1) Point-based methods [39, 41] utilize point-level embeddings for 3D object detection.
-2) Voxel-based methods [55, 10, 59, 1, 56, 12] voxelize the 3D scene and use sparse convolution or transformer for embedding.
-3) Point-voxel-combined methods [40, 42] utilize both embeddings from 1) and 2).
-For camera-based 3D perception, [11, 7, 43, 50, 60, 20] embed 2D features on image plane and project these 2D features into 3D space with estimated depth.
-Recently, research starts to focus on fusion 3D perception [8, 30, 26, 22, 25, 21, 32] with camera images and LiDAR point clouds as inputs.
-These methods mainly focus on integrating embeddings of different modalities and train both the 2D and 3D backbones in an supervised manner.
-As labeling in 3D space is costly, we explore unsupervised 3D representation learning for fusion 3D perception.
+1) Point-based methods (Shi et al., 2019; 2020b) utilize point-level embeddings for 3D object detection.
+2) Voxel-based methods (Yan et al., 2018; Deng et al., 2021; Yin et al., 2021; Bai et al., 2022; Yang et al., 2018; Fan et al., 2021) voxelize the 3D scene and use sparse convolution or transformer for embedding.
+3) Point-voxel-combined methods (Shi et al., 2020a; 2021) utilize both embeddings from 1) and 2).
+For camera-based 3D perception, (Ding et al., 2020; Chen et al., 2016; Simonelli et al., 2019; Wang et al., 2021b; Zhang et al., 2021; Ku et al., 2019) embed 2D features on image plane and project these 2D features into 3D space with estimated depth.
+Recently, research starts to focus on fusion 3D perception (Chen et al., 2017; Liu et al., 2023; Liang et al., 2022; Li et al., 2022b; Liang et al., 2019; Li et al., 2022a; Meyer et al., 2019) with camera images and LiDAR point clouds as inputs.
+These methods focus on integrating embeddings of different modalities and train in an supervised manner.
+There are other works on fusion-based semantic segmentation (Wu et al., 2024) and object tracking (Liu et al., 2024).
+As labeling in 3D space is costly, we explore unsupervised 3D pre-training for fusion 3D perception.
 3D Pre-training.
-Annotating for 3D data is notoriously time- and energy-consuming and the emergence of unsupervised representation learning for 2D image [15, 47, 4, 13, 51, 16] provides a promising way to alleviate the annotation burden.
-Existing works [52, 27, 17, 6, 19, 24, 28, 38, 36, 57, 53, 18, 62] in unsupervised 3D representation learning into scene-level 3D point clouds can be divided into two contrastive-based and masked-and-reconstruction-based paradigms.
-Contrastive-based works [52, 27, 17, 6, 19, 24, 36, 28] propose various ways to build suitable views and conduct contrastive learning to improve the performance in downstream perception task.
-Inspired by [16] in image domain, [57, 53, 18, 62] propose to first mask the input point clouds and pre-train the 3D encoders with a shallow decoder for reconstructing the unmasked inputs.
-[44, 61, 23] are pioneering works to introduce contrastive learning into fusion perception.
+Annotating for 3D data is notoriously time- and energy-consuming and the emergence of unsupervised representation learning for 2D image (He et al., 2020; Tian et al., 2019; Caron et al., 2020; Grill et al., 2020; Wang et al., 2021c; He et al., 2022) provides a promising way to alleviate the annotation burden.
+Existing works  (et al, 2020; Liu et al., 2020; Hou et al., 2021; Chen et al., 2022; Huang et al., 2021; Liang et al., 2021; Liu et al., 2021a; Sautier et al., 2022; Pang et al., 2023; Yang et al., 2023; et al, 2023; Huang et al., 2023; Zhu et al., 2023) in unsupervised 3D representation learning into scene-level 3D point clouds can be divided into two contrastive-based and masked-and-reconstruction-based paradigms.
+Contrastive-based works (et al, 2020; Liu et al., 2020; Hou et al., 2021; Chen et al., 2022; Huang et al., 2021; Liang et al., 2021; Pang et al., 2023; Liu et al., 2021a) propose various ways to build suitable views and conduct contrastive learning to improve the performance in downstream perception task.
+Inspired by (He et al., 2022) in image domain, (Yang et al., 2023; et al, 2023; Huang et al., 2023; Zhu et al., 2023) propose to first mask the input point clouds and pre-train the 3D encoders with a shallow decoder for reconstructing the unmasked inputs.
+(Sun et al., 2023; Zhang et al., 2024; Li et al., 2022c) are pioneering works to introduce contrastive learning into fusion perception.
 They consider camera and LiDAR embeddings as different views of the scene and apply the contrastive loss between the two modalities.
-Inspired by the success of neural field in representing 3D scenes [33, 49] and previous attempts to introduce neural rendering to 3D pre-training for point clouds [54, 18, 62], UniPAD [58] proposes to use a differentiable-rendering decoder for masked-and-reconstruction pre-training and achieves SOTA performance for unsupervised 3D representation learning on fusion 3D perception.
-However, due to the high GPU memory consumption, UniPAD [58] is only able to separately pre-train the image and point cloud encoders and fails to utilize the interaction between modalities during pre-training.
+Inspired by the success of neural field in representing 3D scenes (Mildenhall et al., 2021; Wang et al., 2021a) and previous attempts to introduce neural rendering to 3D pre-training for point clouds (Yan et al., 2023; Huang et al., 2023; Zhu et al., 2023), UniPAD (Yang et al., 2024) proposes to use a differentiable-rendering decoder for masked-and-reconstruction pre-training and achieves SOTA performance for unsupervised 3D representation learning on fusion 3D perception.
+However, due to the high GPU memory consumption, UniPAD (Yang et al., 2024) is only able to separately pre-train the image and point cloud encoders and fails to utilize the interaction between modalities during pre-training.
 In this paper, we explore unsupervised joint pre-training for 2D and 3D backbones via differentiable rendering with Curvature Sampling and Prototype Learning.
-Previous work like SwAV [4] is related to our prototype learning part.
-SwAV [4] explores learnable prototypes to represent different categories in the same modality (image) and uses swapping assignment prediction loss for different views of the same image instance.
-On the contrary, learnable prototypes in CLAP are used to represent part of the 3D scenes and learn the interaction between modalities, which differs from the context in [4].
+Previous work like SwAV (Caron et al., 2020) is related to our prototype learning part.
+SwAV (Caron et al., 2020) explores learnable prototypes to represent different categories in the same modality (image) and uses swapping assignment prediction loss for different views of the same image instance.
+On the contrary, learnable prototypes in CLAP are used to represent part of the 3D scenes and learn the interaction between modalities, which differs from the context in (Caron et al., 2020).
 Thus we propose a Expectation-Maximization training scheme to maximize similarity between prototypes and 3D embeddings.
@@
 In order to jointly pre-train the LiDAR, camera and fusion encoders, we first embed the paired LiDAR point clouds and camera images with $f^{\text{enc}}_{\text{P}}$, $f^{\text{enc}}_{\text{I}}$ and $f^{\text{enc}}_{\text{fusion}}$.
-Then based on the fusion features, CLAP applies differentiable rendering to predict both depth and rgb with the SDF and RGB values of the sampled points along LiDAR/camera rays from $f^{\text{SDF}}$ and $f^{\text{RGB}}$, with which we compute loss against the observed LiDAR point cloud and camera images.
+Then based on the fusion features, CLAP applies differentiable rendering to predict both range and rgb with the SDF and RGB values of the sampled points along LiDAR/camera rays from $f^{\text{SDF}}$ and $f^{\text{RGB}}$, with which we compute loss against the observed LiDAR point cloud and camera images.
 To make joint pre-training feasible, we propose Curvature Sampling to sample informative parts of the 3D scene, as descriped in (a).
@@
 Notations.
-To begin with, we denote the input image set from $N_{\text{cam}}$ cameras as =$\mathcal{I}=\{\mathbf{I}_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N_{\text{cam}}}$ and LiDAR point cloud as $\mathbf{P}\in\mathbb{R}^{N_{p}\times(3+d)}$.
+To begin with, we denote the input image set from $N_{\text{cam}}$ cameras as $\mathcal{I}=\{\mathbf{I}_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N_{\text{cam}}}$ and LiDAR point cloud as $\mathbf{P}\in\mathbb{R}^{N_{p}\times(3+d)}$.
 $H$ and $W$ are the height and width of the images and each pixel on the images has $3$ values for RGB.
-$N_{p}$ is the number of points in the LiDAR point cloud and each of them contains $xyz$-location and $d$ feature channels.
-For example, in NuScenes [3] dataset, $d=2$ represents the intensity and timestamp of each point and there are $N_{\text{cam}}=6$ surrounding cameras on the autonomous vehicle.
+$N_{p}$ is the number of points in the LiDAR point cloud and each of them contains xyzxyz-location and $d$ feature channels.
+For example, in NuScenes (Caesar et al., 2020) dataset, $d=2$ represents the intensity and timestamp of each point and there are $N_{\text{cam}}=6$ surrounding cameras on the autonomous vehicle.
 For each pair of camera image and LiDAR point cloud, we have the transformation matrix $\mathbf{T}_{n}\in\mathbb{R}^{3\times 4}$ indicating the projection between the camera plane and LiDAR coordinate, where $n=1,2,...,N_{\text{cam}}$.
@@
 $\hat{D}$, $\hat{H}$ and $\hat{W}$ are spatial resolutions of the embedded features and $\hat{d}_{\text{P}}$ is number of feature channels after encoding.
-Then for camera images $\mathcal{I}$, $f^{\text{enc}}_{\text{I}}$ encodes them with swin transformer [29] and uses $\mathcal{T}=\{\mathbf{T}_{n}\}_{n=1}^{N_{\text{cam}}}$ to project the 2D features to 3D space.
+Then for camera images $\mathcal{I}$, $f^{\text{enc}}_{\text{I}}$ encodes them with swin transformer (Liu et al., 2021b) and uses $\mathcal{T}=\{\mathbf{T}_{n}\}_{n=1}^{N_{\text{cam}}}$ to project the 2D features to 3D space.
 $\hat{\mathbf{I}}=f^{\text{enc}}_{\text{I}}(\mathcal{I},\mathcal{T},\mathbf{P}),$ (2)
 where $\hat{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{I}}}$ is the embedded 3D features for surrounding camera images with the similar dimensions as $\hat{\mathbf{P}}$ except for $\hat{d}_{\text{I}}$ feature channels.
-The projection of 2D features to 3D space is similar to [30]: we transform LiDAR points back to image planes with $\mathcal{T}$ and use the projected depths to project the 2D features to 3D space.
+The projection of 2D features to 3D space is similar to (Liu et al., 2023): we transform LiDAR points back to image planes with $\mathcal{T}$ and use the projected ranges from LiDAR to project the 2D features to 3D space.
 With $\hat{\mathbf{P}}$ and $\hat{\mathbf{I}}$, we further concatenate them along feature dimension and apply the fusion encoder $f^{\text{enc}}_{\text{fusion}}$ to get the fusion feature $\hat{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$ with $\hat{d}_{\text{F}}$ feature dimensions,
 $\hat{\mathbf{F}}=f^{\text{enc}}_{\text{fusion}}([\mathbf{P},\mathbf{I}]).$ (3)
-Loss Function.
-To guide $f^{\text{enc}}_{\text{P}}$, $f^{\text{enc}}_{\text{I}}$ and $f^{\text{enc}}_{\text{fusion}}$ to learn good representations in an unsupervised manner, CLAP first embed the fusion features $\hat{\mathbf{F}}$ with a shallow 3D convolution network $f^{\text{3D}}$ to get $\tilde{\mathbf{F}}=f^{\text{3D}}(\hat{\mathbf{F}})$ and we have $\tilde{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$.
-Then a rendering loss $\mathcal{L}_{\text{rend}}$ is applied on $\tilde{\mathbf{F}}$ for masked-reconstruction on both point clouds $\mathbf{P}$ and images $\mathcal{I}$.
-Furthermore, a prototype learning scheme $\mathcal{L}_{\text{proto}}$ is utilized in order to bridge the two modalities and incorporate interaction of image semantics and LiDAR geometry into pre-training.
-The overall loss function is as below:
-$\mathcal{L}=\omega_{\text{r}}\times\mathcal{L}_{\text{rend}}(\mathbf{P},\tilde{\mathbf{F}},\mathcal{I})+\omega_{\text{proto}}\times\mathcal{L}_{\text{proto}}(\hat{\mathbf{P}},\hat{\mathbf{I}}),$ (4)
+[…]
+[…]
 with $\omega_{\text{r}}$ and $\omega_{\text{proto}}$ as weighting parameters to balance the losses.
 3.2 Neural Field and Differentiable Rendering
-Inspired by the success of UniPAD [58], CLAP applies a differentiable rendering decoder with neural field to conduct the masked-and-reconstruction pre-training.
-Different from [58], an additional surface signed distance field loss is applied to better optimize the scene geometry.
-Here we first introduce Neural Field, which is the basis for camera images and point clouds rendering, and then discuss the differentiable rendering process on depth and RGB values.
+Inspired by the success of UniPAD (Yang et al., 2024), CLAP applies a differentiable rendering decoder with neural field to conduct the masked-and-reconstruction pre-training.
+Different from (Yang et al., 2024), an additional surface signed distance field loss is applied to better optimize the scene geometry.
+Here we first introduce Neural Field, which is the basis for camera images and point clouds rendering, and then discuss the differentiable rendering process on range and RGB values.
 Neural Field.
 Given a specific point $\mathbf{p}=[x,y,z]\in\mathbb{R}^{3}$ in the 3D space, the feature $\mathbf{f}_{\text{p}}\in\mathbb{R}^{\hat{d}_{\text{F}}}$ at $\mathbf{p}$ is queried from the fusion 3D embedding $\tilde{\mathbf{F}}$ by trilinear interpolation
-$\mathbf{f}_{\text{p}}=f^{\text{tri}}(\mathbf{p},\hat{\mathbf{P}}_{t}),$ (5)
-where $f^{\text{tri}}$ is an built-in module implemented in Pytorch [37].
-Taking the concatenation of location $\mathbf{p}$ and queried feature $\mathbf{f}_{\text{p}}$ as inputs, we predict the signed distance value $s\in\mathbb{R}$ [5, 31] and color value $c\in\mathbb{R}^{3}$ [49] at $\mathbf{p}$ with $f^{\text{SDF}}$ and $f^{\text{RGB}}$.
+[…]
+where $f^{\text{tri}}$ is an built-in module implemented in Pytorch (Paszke et al., 2019).
+Taking the concatenation of location $\mathbf{p}$ and queried feature $\mathbf{f}_{\text{p}}$ as inputs, we predict the signed distance value $s\in\mathbb{R}$ (Chan & Zhu, 2005; Malladi et al., 1995) and color value $c\in\mathbb{R}^{3}$ (Wang et al., 2021a) at $\mathbf{p}$ with $f^{\text{SDF}}$ and $f^{\text{RGB}}$.
+$f^{\text{RGB}}$ and $f^{\text{SDF}}$ are parameterized by Multi-layer Perceptron.
 $s=f^{\text{SDF}}([\mathbf{p},\mathbf{f}])\ \ ,\ \ c=f^{\text{RGB}}([\mathbf{p},\mathbf{f}]),$ (6)
-where $f^{\text{RGB}}$ and $f^{\text{SDF}}$ are parameterized by Multi-layer Perceptron.
 Differentiable Rendering.
-Similar to [33, 49], we first sample $N_{\text{L}}$ or $N_{\text{C}}$ rays at the LiDAR or camera sensor origin $\mathbf{o}$, each of which is described by its normalized direction $\mathbf{d}$ and $\mathbf{o}$.
-Next, we sample $N_{\text{ray}}$ points following [49] along each ray.
+Similar to (Mildenhall et al., 2021; Wang et al., 2021a), we first sample $N_{\text{L}}$ or $N_{\text{C}}$ rays at the LiDAR or camera sensor origin $\mathbf{o}$, each of which is described by its normalized direction $\mathbf{d}$ and $\mathbf{o}$.
+Next, we sample $N_{\text{ray}}$ points following (Wang et al., 2021a) along each ray.
 Here each point along the ray can be interpreted by $\mathbf{p}=\mathbf{o}+r\mathbf{d}$, where $r$ is the range from the sensor origin to the point $\mathbf{p}$.
 Thus the sampled point set can be annotated by $\{\mathbf{p}_{n}=\mathbf{o}+r_{n}\mathbf{d}\}_{n=1}^{N_{\text{ray}}}$ and we predict the estimated signed distance value $s_{n}$ and color value $c_{n}$ for them with $f^{\text{RGB}}$ and $f^{\text{SDF}}$.
-Following [49], we estimate the occupancy value $\alpha_{n}$ for each sampled point,
+Following (Wang et al., 2021a), we estimate the occupancy value $\alpha_{n}$ for each sampled point,
 $\alpha_{n}=\max{(\frac{\Phi_{h}(s_{n})-\Phi_{h}(s_{n+1})}{\Phi_{h}(s_{n})},0)}.$ (7)
-Here $\Phi_{h}(x)=(1+e^{-hx})^{-1}$ stands for the sigmoid function paired with a learnable scalar $h$. After that, we predict the accumulated transmittance $t_{n}$ similar to [49]
+Here $\Phi_{h}(x)=(1+e^{-hx})^{-1}$ stands for the sigmoid function paired with a learnable scalar $h$. After that, we predict the accumulated transmittance $t_{n}$ similar to (Wang et al., 2021a)
 $t_{n}=\prod^{n-1}_{i=1}(1-\alpha_{i}).$ (8)
-Based on $t_{n}$, we compute an unbiased and occlusion-aware weight $w_{n}=t_{n}\alpha_{n}$ [49] and integrate all samples along the ray to predict the range $\tilde{r}$ or color $\tilde{c}$ along this ray,
+Based on $t_{n}$, we compute an unbiased and occlusion-aware weight $w_{n}=t_{n}\alpha_{n}$ (Wang et al., 2021a) and integrate all samples along the ray to predict the range $\tilde{r}$ or color $\tilde{c}$ along this ray,
 $\tilde{r}=\sum_{n=1}^{N_{\text{ray}}}w_{n}*r_{n}\ \ ,\ \ \tilde{c}=\sum_{n=1}^{N_{\text{ray}}}w_{n}*c_{n}$ (9)
-For the observed LiDAR points, it is evident that signed distance value at those points are 0.
-The final loss function for differentiable rendering is a combination of L-1 loss on range and color predictions and the surface SDF loss.
+For the observed LiDAR points, it is evident that signed distance value should be 0.
+The loss function for differentiable rendering is a combination of L-1 loss on surface SDF, range and color predictions.
 $\mathcal{L}_{\text{rend}}=\frac{1}{N_{\text{L}}}\sum_{i=1}^{N_{\text{L}}}(|r_{i}-\tilde{r}_{i}|+\omega_{\text{sur}}|s_{i}|)+\frac{\omega_{\text{C}}}{3\cdot N_{\text{C}}}\sum_{i=1}^{N_{\text{C}}}\sum_{j=1}^{3}|c_{i}^{j}-\tilde{c}_{i}^{j}|,$ (10)
@@
 $c_{i}^{j}$ is the value of $j^{th}$ channel in the image pixel, where $j=\{1,2,3\}$ corresponds to RGB channels.
-Figure 3: Inspiration of Curvature Sampling.
-It can be found that areas with low curvature tend to be less informative, like road plane, while those with high curvature provide more information (objects like cars).
 3.3 Curvature Sampling
 In order to make joint unsupervised representation learning feasible, we have to make $N_{\text{L}}\ll N_{\text{P}}$ and $N_{\text{C}}\ll H\cdot W\cdot N_{\text{cam}}$.
-Intuitively, uniform sampling with depth can be used, same as “Memory-friendly Ray Sampling” in [58].
+Intuitively, uniform sampling with range can be used,
+Figure 3: Inspiration of Curvature Sampling.
+same as “Memory-friendly Ray Sampling” in (Yang et al., 2024).
 But due to the relatively small sample number compared to the raw inputs ($\sim\frac{1}{100}$), this sampling method brings little improvement against separate pre-training, which is contradictory to our motivation.
@@
 Therefore, we propose the Curvature Sampling for effective sampling.
-For each point $\mathbf{p}$ in the LiDAR point cloud $\mathbf{P}$, we first estimate the surface normal by deriving the signed distance function with respect to $\mathbf{p}$
+In order to make it more intuitive, we show camera image here and the curvature estimation is actually computed in 3D with the SDF function.
+For each point $\mathbf{p}$ in the LiDAR point cloud $\mathbf{P}$, we first estimate surface normal by deriving the signed distance function with respect to $\mathbf{p}$
 $\mathbf{n}=\frac{\delta f^{\text{SDF}}([\mathbf{p},\mathbf{f}])}{\delta\mathbf{p}},$ (11)
-where $\mathbf{n}\in\mathbb{R}^{3}$ is the predicted normal.
-Then we normalized $\mathbf{n}$ to get the direction of the normal $\tilde{\mathbf{n}}=\frac{\mathbf{n}}{||\mathbf{n}||_{2}}$.
-Here $||\cdot||_{2}$ is the L-2 norm.
-Next we apply the differential operator on $\tilde{\mathbf{n}}$ with respect to $\mathbf{p}$ and get $\mathbf{c}\in\mathbb{R}^{3}$.
+[…]
 $\mathbf{c}=\frac{\delta\tilde{\mathbf{n}}}{\delta\mathbf{p}}.$ (12)
-Then we estimate the geodesic curvature [48] of each point $\mathbf{p}_{n}$ in LiDAR point cloud by computing the norm of $\mathbf{c}_{n}$ and use it as the sampling weights $\omega_{n}$, that is $\omega_{n}=||\mathbf{c}_{n}||_{2}$.
-With $\omega_{n}$, we sample $N_{\text{L}}$ points with a Multinomial Sampler implemented in PyTorch [37] for differentiable rendering.
-For pixels on image plane, we project the LiDAR point cloud back to image planes with $\mathcal{T}$, assign $\omega_{n}$ of each point to the projected pixel and apply a gaussian blur kernel of size $K_{\text{gaus}}$ to densify the weights, with which we sample $N_{\text{C}}$ pixels for $\mathcal{L}_{\text{rend}}$.
-As the estimation of curvature here is noisy especially in the first few epochs, we set up a warm-up stage with $N_{\text{warmup}}$ epochs with uniform sampling and after $N_{\text{warmup}}$ epochs, Curvature Sampling is utilized.
+[…]
 3.4 Prototype Learning
@@
 Expectation-Maximization.
-In order to guide the learnable prototypes to represent parts of the environment, we propose an Expectation-Maximization [35] training scheme to optimize the prototypes.
+In order to guide the learnable prototypes to represent parts of the environment, we propose an Expectation-Maximization (Moon, 1996) training scheme to optimize the prototypes.
 We first project the LiDAR embeddings $\hat{\mathbf{P}}$ and camera embeddings $\hat{\mathbf{I}}$ separately with two projection heads $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ to the same dimension as $\mathbf{K}$
-$\dot{\mathbf{P}}=f^{\text{proj}}_{\text{P}}(\hat{\mathbf{P}})\ \ ,\ \ \dot{\mathbf{I}}=f^{\text{proj}}_{\text{I}}(\hat{\mathbf{I}}),$ (13)
+[…]
 where $f^{\text{proj}}_{\text{P}}$ and $f^{\text{proj}}_{\text{I}}$ are parameterized by Multi-Layer Perceptron and $\dot{\mathbf{P}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times d_{\text{K}}}$, $\dot{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times d_{\text{K}}}$.
@@
 In the Expectation step, we compute the probability $\hat{\mathbf{S}}_{\text{P/I}}$ that each prototype is assigned to each embeddings by applying a softmax operation on $\mathbf{S}_{\text{P/I}}$.
-Then for Maximization step, we expect to maximize the probability of the assignment between one prototype to one specific part of the scene and this is the same as minimizing the entropy of the similarity matrix.
-Hence, the EM loss is computed as,
-$\mathcal{L}_{\text{EM}}=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\hat{\mathbf{S}}_{\text{P}}^{n,m}log\hat{\mathbf{S}}_{\text{P}}^{n,m}+\hat{\mathbf{S}}_{\text{I}}^{n,m}log\hat{\mathbf{S}}_{\text{I}}^{n,m}\}$ (15)
+Then for Maximization step, we expect to maximize the probability of the assignment between one prototype to one specific part of the scene and this is equal to minimize the entropy of the similarity matrix.
+Thus, the EM loss is computed as
+$\mathcal{L}_{\text{EM}}=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\hat{\mathbf{S}}_{\text{P}}^{n,m}log\hat{\mathbf{S}}_{\text{P}}^{n,m}+\hat{\mathbf{S}}_{\text{I}}^{n,m}log\hat{\mathbf{S}}_{\text{I}}^{n,m}\}.$ (15)
 Swapping Prototype Prediction.
-To further explore interaction between modalities, we detach $\mathbf{S}_{\text{P/I}}$ and apply sinkhorn algorithm [9] to make them approximate to double stochastic matrix in $N_{\text{sink}}$ iterations.
+To further explore interaction between modalities, we detach $\mathbf{S}_{\text{P/I}}$ and apply sinkhorn algorithm (Cuturi, 2013) to approximate them to double stochastic matrix in $N_{\text{sink}}$ iterations.
 We denote the updated matrix as codes $\mathbf{Q}_{\text{P/I}}\in\mathbb{R}^{N_{\text{3D}}\times N_{\text{K}}}$.
-The swapping prototype prediction loss is formulated with a temperature parameter $\tau$, inspired by [4],
+The swapping prototype prediction loss is computed with a temperature parameter $\tau$, inspired by (Caron et al., 2020),
 $\begin{split}\mathcal{L}_{\text{SwAV}}&=-\frac{1}{N_{\text{3D}}N_{\text{K}}}\sum_{n=1}^{N_{\text{3D}}}\sum_{m=1}^{N_{\text{K}}}\{\mathbf{Q}_{\text{I}}^{n,m}\log\frac{exp(\mathbf{S}_{\text{P}}^{n,m})/\tau}{\sum_{k=1}^{N_{\text{K}}}exp(\mathbf{S}_{\text{P}}^{n,k})/\tau}\\ &+\mathbf{Q}_{\text{P}}^{n,m}\log\frac{exp(\mathbf{S}_{\text{I}}^{n,m})/\tau}{\sum_{k=1}^{N_{\text{K}}}exp(\mathbf{S}_{\text{I}}^{n,k})/\tau}\}.\end{split}$ (16)
 Gram Matrix Minimization.
-When training the randomly initialized prototypes, the network might learn a short cut with all prototypes being the same [4], which is called collapse.
+When training the randomly initialized prototypes, the network might learn a short cut with all prototypes being the same (Caron et al., 2020), which is called collapse.
 To avoid this, we estimate similarity between prototypes by the gram matrix $\mathbf{G}=\mathbf{K}\mathbf{K}^{\intercal}$ of prototypes $\mathbf{K}$, the dimension of which is $\mathbf{G}\in\mathbb{R}^{N_{\text{K}}\times N_{\text{K}}}$.
@@
 $\mathcal{L}_{\text{proto}}=\omega_{\text{SwAV}}\mathcal{L}_{\text{SwAV}}+\omega_{\text{EM}}\mathcal{L}_{\text{EM}}+\omega_{\text{GMM}}\mathcal{L}_{\text{GMM}}.$ (18)
+4 Experiments
+Unsupervised 3D representation learning for fusion perception aims to pre-train both LiDAR and camera encoders and initialize downstream models with the pre-trained weights to gain performance improvement in downstream tasks.
+In this section, we design extensive experiments on the popular autonomous driving dataset NuScenes (Caesar et al., 2020) and Waymo (Sun et al., 2020) to demonstrate the effectiveness of CLAP.
+To begin with, we describe experiment setups in Section 4.1.
+Next, we show and analyze main results in Section 4.2.
+Finally, we provide ablation study and visualizations separately in Section 4.3 and 4.5.
+4.1 Settings
+Datasets.
+We use the popular autonomous driving dataset NuScenes (Caesar et al., 2020) and Waymo (Sun et al., 2020) to evaluate the performance of CLAP.
+NuScenes (Caesar et al., 2020) uses one roof LiDAR and six surrounding cameras to collect data.
+The LiDAR is a 32-beam Velodyne and collecting frequency is 20Hz.
+The frequency of camera capturing is 12Hz.
+(Caesar et al., 2020) conducts the synchronization and provides paired data of LiDAR point cloud and camera images.
+The whole NuScenes dataset contains 1000 scenes collected in Boston and Singapore.
+Each scene lasts for around 20 seconds and there are a total of 5.5 hours data.
+Following convention practice from (Caesar et al., 2020; Team, 2020), we divide the whole dataset into training set with 850 scenes and validation set with 150 scenes.
+Waymo (Sun et al., 2020) uses one top 64-beam LiDAR, 4 corner LiDARs and 5 surrounding cameras to collect point clouds and camera images.
+Following the same practice in (Sun et al., 2020; Team, 2020), the collected 1000 scenes are split into training set (798 scenes) and validation set (202 scenes).
+All pre-trainings are conducted on the training set without labels and we conduct downstream 3D object detection training in few-shot setting via uniform sampling of the training data (NuScenes $5\%$ and Waymo $1\%$).
+Downstream 3D Object Detectors.
+For NuScenes, we select the SOTA 3D object detector called BEVFusion (Liu et al., 2023) and for Waymo, we use CenterPoint (Yin et al., 2021).
+Both BEVFusion and CenterPoint are implemented in the popular code repository for autonomous driving perception called OpenPCDet (Team, 2020).
+For evaluation metrics, we use average precisions of various categories (APs), mean average precision (mAP) and NuScenes Detection Score (NDS) (Caesar et al., 2020) for NuScenes and mAP (mean accurate precisions) and mAPH (mean accurate precisions with headings) at different difficulty levels (Level-1 and 2) for Waymo (Sun et al., 2020).
+We follow a similar setting in (He et al., 2019; et al, 2023) to gradually increase training iterations of the from-scratch model until convergence is observed.
+Here convergence means further increasing training iterations will not improve the performance.
+Then the number of training iterations is fixed for fine-tuning pre-trained models.
+This setting avoids the case that pre-training only accelerates convergence and makes sure that pre-training indeed improve the performance of downstream models, that is improving the sample efficiency of the downstream task.
+Baseline Pre-training Method for Fusion Perception.
+We incorporate three kinds of pre-training baseline methods: 1) an occupancy estimation method called ALSO (Boulch et al., 2023), 2) occupancy masked autoencoder called Occupancy-MAE (Min et al., 2023), 3) multi-modality methods including SLidR (Sautier et al., 2022), PPKT (Liu et al., 2021a) and UniPAD (Yang et al., 2024).
+We use the official implementations to pre-train the backbones with same setting as CLAP.
+Table 1: Results for fine-tuning on 5% of training set in NuScenes.
+“Init.” means the way to initialize models.
+We guarantee convergence of from-scratch model and fix the training iteration for all fine-tuning experiments.
+We provide mAP and NDS as an evaluation of the overall performance of different models and highlight the best mAP and NDS with bold font.
+We also indicate the performance improvement by green color.
+“C.V.”, “Mot.”, “Bic.”, “Ped.” and “T.C.” are abbreviations for Construction Vehicle, Motorcycle, Bicycle, Pedestrian and Traffic Cone.
+Results are in %.
 Init. mAP NDS Car Truck C.V. Bus Trailer Barrier Mot. Bic. Ped. T.C.
@@
 CLAP 51.17 +2.48 57.04 +1.76 79.56 48.43 18.84 56.34 23.98 60.60 48.87 34.11 78.08 62.87
-Table 1: Results for fine-tuning on 5% of training set in NuScenes [3].
-“Init.” means the way to initialize models.
-To avoid the case the performance difference is brought by accelerating convergence, we first increase the training iteration of training-from-scratch model (BEVFusion [30]) until we observe convergence, which leads to the “Rand.” row.
-Then we pre-train the backbones with ALSO [2], OCC-MAE [34], SLidR [38], PPKT [28], UniPAD [58] and CLAP.
-We initialize BEVFusion with the pre-trained weights and fine-tune the model with the same number iterations as “Rand.”.
-We provide mAP and NDS as an evaluation of the overall performance of different models and highlight the best mAP and NDS with bold font.
-We also indicate the performance improvement by green color.
-“C.V.”, “Mot.”, “Bic.”, “Ped.” and “T.C.” are abbreviations for Construction Vehicle, Motorcycle, Bicycle, Pedestrian and Traffic Cone.
-All the results here are in %.
-4 Experiments
-Unsupervised 3D representation learning for fusion perception aims to pre-train both LiDAR and camera encoders and initialize downstream models with the pre-trained weights to gain performance improvement in downstream tasks.
-In this section, we design extensive experiments on the popular autonomous driving dataset NuScenes [3] and Waymo [45] to demonstrate the effectiveness of CLAP.
-To begin with, we describe experiment setups in Section 4.1.
-Next, we show and analyze main results in Section 4.2.
-Finally, we provide ablation study and visualizations separately in Section 4.3 and 4.4.
-4.1 Settings
-Datasets.
-We use the popular autonomous driving dataset NuScenes [3] and Waymo [45] to evaluate the performance of CLAP.
-NuScenes [3] uses one roof LiDAR and six surrounding cameras to collect data.
-The LiDAR is a 32-beam Velodyne and collecting frequency is 20Hz.
-The frequency of camera capturing is 12Hz.
-[3] conducts the synchronization and provides paired data of LiDAR point cloud and camera images.
-The whole NuScenes dataset contains 1000 scenes collected in Boston and Singapore.
-Each scene lasts for around 20 seconds and there are a total of 5.5 hours data.
-Following convention practice from [3, 46], we divide the whole dataset into training set with 850 scenes and validation set with 150 scenes.
-Waymo [45] uses one top 64-beam LiDAR, 4 corner LiDARs and 5 surrounding cameras to collect point clouds and camera images.
-Following the same practice in [45, 46], the collected 1000 scenes are split into training set (798 scenes) and validation set (202 scenes).
-All pre-trainings are conducted on the training set without labels and we conduct downstream 3D object detection training in few-shot setting via uniform sampling of the training data (NuScenes [3] $5\%$ and Waymo [45] $1\%$).
-Downstream 3D Object Detectors.
-For NuScenes [3], we select the SOTA 3D object detector called BEVFusion [30] and for Waymo [45], we use CenterPoint [59].
-Both BEVFusion and CenterPoint are implemented in the popular code repository for autonomous driving perception called OpenPCDet [46].
-For evaluation metrics, we use average precisions of various categories (APs), mean average precision (mAP) and NuScenes Detection Score (NDS) [3] for NuScenes [3] and mAP (mean accurate precisions) and mAPH (mean accurate precisions with headings) at different difficulty levels (Level-1 and 2) for Waymo [45].
-We follow a similar setting in [14, 53] to gradually increase training iterations of the from-scratch model until convergence is observed.
-Here convergence means further increasing training iterations will not improve the performance.
-Then the number of training iterations is fixed for fine-tuning pre-trained models.
-This setting avoids the case that pre-training only accelerates convergence and makes sure that pre-training indeed improve the performance of downstream models, that is improving the sample efficiency of the downstream task.
-Baseline Pre-training Method for Fusion Perception.
-We incorporate three kinds of pre-training baseline methods: 1) an occupancy estimation method called ALSO [2], 2) occupancy masked autoencoder called Occupancy-MAE [34], 3) multi-modality methods including SLidR [38], PPKT [28] and UniPAD [58].
-We use the official implementations to pre-train the backbones with same setting as CLAP.
 Implementation Details of CLAP.
@@
 The loss weighting parameters are implemented as $\omega_{\text{r}}=2.0$, $\omega_{\text{proto}}=1.0$, $\omega_{\text{sur}}=0.05$, $\omega_{\text{C}}=0.05$, $\omega_{\text{SwAV}}=1.0$, $\omega_{\text{EM}}=0.1$ and $\omega_{\text{GMM}}=0.1$.
-We use torch.auto_grad() [37] to implement the derivatives in the curvature estimation.
-For more details about pre-training and fine-tuning, please refer to Appendix 1.
+We use torch.auto_grad() (Paszke et al., 2019) to implement the derivatives in the curvature estimation.
+More details about pre-training and fine-tuning are provided in Appendix A.
 4.2 Main Results
 NuScenes Results.
-As shown in Table 1, CLAP achieves 2.48% mAP improvement over randomly initialization at convergence, which is $100\%$ more improvement for mAP than SOTA unsupervised 3D representation method UniPAD [58] and the best among all initialization methods.
-For NDS metric, UniPAD [58] only achieves comparable performance and PPKT [28] makes a gain of 0.57% while CLAP surpasses the train-from-scratch model by 1.76% .
+As shown in Table 1, CLAP achieves 2.48% mAP improvement over randomly initialization at convergence, which is $100\%$ more improvement for mAP than SOTA unsupervised 3D representation method UniPAD (Yang et al., 2024) and the best among all initialization methods.
+For NDS metric, UniPAD (Yang et al., 2024) only achieves comparable performance and PPKT (Liu et al., 2021a) makes a gain of 0.57% while CLAP surpasses the train-from-scratch model by 1.76% .
 When it turns to different categories, CLAP generally benefit the performance of all the categories and for Construction Vehicle, Bus, Barrier, Motorcycle and Bicycle, the improvement over random initialization are more than 2% AP.
+Waymo Results.
+In Table 4.2, we provide the average performance difference of mAP and mAPH over different difficulty levels.
+It can be found that CLAP achieves the best performance at convergence.
+Meanwhile, the performance gain brought by CLAP is approximately two times as the best (OCC-MAE (Min et al., 2023)) of previous pre-training methods.
+This demonstrates the effectiveness and generalization ability of CLAP.
+Table 2: Results for Waymo (Sun et al., 2020).
 Init. Level-1 Level-2 $\bar{\Delta}$
@@
 CLAP 62.87 59.88 56.88 54.16 $\color[rgb]{0.1328125,0.546875,0.1328125}{\mathbf{+1.28}}$
-Table 2: Results for fine-tuning on $1\%$ of training set in Waymo [45].
-We first increase training iterations of from-scratch model and observe convergence (Rand.).
-Then we fix the training schedule to fine-tune the pre-training models.
-We provide an average difference $\bar{\Delta}$ to show the performance gain.
-All results are in %.
-Waymo Results.
-In Table 2, we provide the average performance difference of mAP and mAPH over different difficulty levels.
-It can be found that CLAP achieves the best performance at convergence.
-Meanwhile, the performance gain brought by CLAP is approximately two times as the best (OCC-MAE [34]) of previous pre-training methods.
-This demonstrates the effectiveness and generalization ability of CLAP.
-Potential Scaling Property.
-As we are not able to scale up the pre-training dataset at current stage, we explore potential scaling property by gradually decreasing the sample numbers (2.5%, 1% and 0.5%) for fine-tuning on NuScenes, which increases the ratio between pre-training data and fine-tuning data.
-The results are shown in Table 3.
-It can be found that as the ratio between pre-training data and fine-tuning data gets larger, the performance improvement by CLAP increases and CLAP provides a gain up to 7.22% mAP and 4.71% NDS with 0.5% fine-tuning data.
-These results show that CLAP is promising in scaling property and in the future, if we can scale up the pre-training dataset, CLAP might further improve current SOTA performance.
-Init. Fine-tune Data mAP NDS
+Table 3: Results on scaling property.
+Init. F.T. Data mAP NDS
 Random 5% 48.69 55.28
@@
 CLAP 23.71 +7.22 27.32 +4.71
-Table 3: Results on potential scaling property of CLAP .
-“Fine-tune Data” stands for the sample number of training set in NuScenes [3].
-Results for random initialization are at convergence and fine-tuning with CLAP uses the same numbers of iterations.
-Results are mAP and NDS in %.
-Figure 4: Visualization of curvature estimation.
-The color of the points change from blue to red, indicating lower sample weight / curvature to higher ones.
-We highlight those correct regions with orange boxes and those noisy estimation with green ones.
-Some of the correct estimated regions are further zoom-in for better understanding.
-Best view in color.
-4.3 Visualizations
-We use the pre-trained model to estimate the curvature of LiDAR point clouds and use heatmap color to indicate the weight computed in Section 3.3.
-The visualization results are shown in Figure 4.
-We use orange boxes to highlight those regions with relatively correct estimation and green ones for those with noisy estimation.
-Furthermore, we zoom in some orange boxes.
-It can be found that though some noise exists (because pre-training is conducted in unsupervised manner), CLAP is able to predict high weights for those highly informative region for sampling and meanwhile assign lower weights to most of the background, which makes joint pre-training feasible.
+Table 4: Ablation Study. The second line is joint pre-training with sampling method from UniPAD.
 Joint Pre-train Cur. Sam. Proto. Learning mAP
@@
 51.17
-[…]
+Figure 4: Visualization of curvature estimation.
+Potential Scaling Property.
+As we are not able to scale up the pre-training dataset at current stage, we explore potential scaling property by gradually decreasing the sample numbers (2.5%, 1% and 0.5%) for fine-tuning on NuScenes, which increases the ratio between pre-training data and fine-tuning data.
+The results are shown in Table 4.2 (F.T.
+is fine-tuning).
+It can be found that as the ratio between pre-training data and fine-tuning data gets larger, the performance improvement by CLAP increases and CLAP provides a gain up to 7.22% mAP and 4.71% NDS with 0.5% fine-tuning data.
+These results show that CLAP is promising in scaling property and in the future, if we can scale up the pre-training dataset, CLAP might further improve current SOTA performance.
+4.3 Visualizations on Curvature Estimation
+We use CLAP to estimate the curvature of point clouds and use heatmap color to indicate the weight computed in Section 3.3.
+As shown in Figure 4 (orange boxes for those regions with relatively correct estimation and green ones for those with noisy estimation), it can be found that though some noise exists (because pre-training is conducted in unsupervised manner), CLAP is able to predict high weights for those highly informative region for sampling and meanwhile assign lower weights to most of the background, which makes joint pre-training feasible.
+4.4 Visualizations on Prototype Learning
+We use the model pre-trained by CLAP to infer the 3D features and assign prototypes to different LiDAR points in the 3D space.
+Then we use different random colors to indicate different prototypes and visualize them, as shown in Figure 5.
+If can be found that the background road plane inside the same frame is generally assigned to the same prototype.
+And foreground vehicles are assigned to another prototype.
+This demonstrates that the proposed prototype learning scheme actually learns to represent parts of the scenes with prototypes in an unsupervised manner.
+However, as our pre-training does not incorporate any label, it can also be found that the prototype assignment has some noise, for example some of the road plane points are assigned to other prototypes.
+Figure 5: Visualization of the prototype learning results. Different color indicates different prototype assignments.
 4.5 Ablation Study
 We conduct ablation study to evaluate the effectiveness of different components.
-Results are in Table 4.
-The first line is seperate pre-training with UniPAD [58].
-The second one is jointly pre-training using UniPAD [58] and uniform sampling guided by depth (“Memory-friendly Ray Sampling” in [58]) to address the GPU memory limitation.
+Results are in Table 4.2.
+The first line is seperate pre-training with UniPAD (Yang et al., 2024).
+The second one is jointly pre-training using UniPAD (Yang et al., 2024) and uniform sampling guided by range (“Memory-friendly Ray Sampling” in (Yang et al., 2024)) to address the GPU memory limitation.
 It can be found that using simple sampling method does not bring improvement over separate pre-training.
-Then we add Curvature Sampling in third line and found that Curvature Sampling enables more effective sampling and improves the performance over uniform sampling guided by depth and separate pre-training.
+Then we add Curvature Sampling in third line and found that Curvature Sampling enables more effective sampling and improves the performance over uniform sampling guided by range and separate pre-training.
 Finally, the Prototype Learning scheme (fourth line) learns a common feature space for segments of 3D scenes and introduces interaction of LiDAR and camera encoders, which achieves the best performance.
@@
 CLAP uses Curvature Sampling to sample more informative parts and learnable prototypes to represent parts of 3D scenes, optimized with an Expectation-Maximization approach.
-To further make use of the interplay between LiDAR geometry and image semantics, a swapping prediction loss is used.
-Finally, a Gram Matrix Regularization loss is proposed to avoid collapse.
+To explore the interplay between LiDAR geometry and image semantics, a swapping prediction loss is used with a regularization loss to avoid collapse.
 Experiment results demonstrate that CLAP is superior in unsupervised 3D representation learning and has the potential to scale up.
+Acknowledgments
+This paper is supported by the General Research Fund of Hong Kong No.17209324 and 17200622, and the Global Industrial Technology Cooperation Center (GITCC) through a grant agreement with the Korea Institute for Advancement of Technology (KIAT), project number P0028922.
+Ethics statement
+We further discuss potential negative social impact of CLAP in this section.
+Job Displacement.
+Automation of tasks that require 3D perception, such as autonomous vehicles or robotics, might lead to job losses in sectors like transportation, warehousing, and manufacturing.
+While automation can create new jobs, the transition period can be challenging for those whose jobs are automated away.
+Security Risks. The deployment of autonomous systems could introduce new vulnerabilities. Any failures in autonomous driving or robotics could lead to accidents or be exploited maliciously.
+Accessibility and Inequality.
+The benefits of advanced 3D perception technologies might not be evenly distributed across society.
+Wealthier regions or organizations may have earlier and better access to these technologies, potentially widening the gap between different socioeconomic groups.
+Reproducibility Statement
+We provide method details in Section 4 and implementation details in Appendix A, which are enough for reproducing the experiments. Besides, we will release the code and models.
+Appendix A More Implementation Details
+Pre-training.
+We use a learning rate of 0.00005 with a cosine learning schedule for pre-training and use mask augmentation for CLAP with a masking rate of 0.9.
+All the pre-trainings are conducted on 8-H100 clusters for similar time ($\sim$ 45 mins / epoch).
+As pre-trained image backbones are broadly used for image feature extraction, the implementation of Yang et al.
+(2024) also use a pre-trained image backbone from Contributors (2020) to initiate their pre-training and the training set for this backbone does not involve any data from NuScenes Caesar et al.
+(2020) and Waymo Sun et al.
+(2020).
+We adopt this practice for training-from-scratch model and CLAP.
+Downstream Training.
+We follow the common practice in OpenPCDet Team (2020) and only change the training iterations in order to observe the convergence of train-from-scratch models, which avoids the case that pre-training only accelerate convergence and make sure that pre-training indeed improve the performance of downstream models.
+The training epoch is 108 for BEVfusion Liu et al.
+(2023) with $5\%$ of NuScenes Caesar et al.
+(2020) training data and 252 for CenterPoint Yin et al.
+(2021) with $1\%$ of Waymo Sun et al.
+(2020) training data.
+Appendix B More Experiment Results
+B.1 Semantic Segmentation
+[…]
+B.2 Repeated Evaluation
+We use the same fixed random seed for all experiments in the main paper for reproducibility.
+As repeated evaluation can further reveal the training robustness, we repeat fine-tunings on NuScenes for 5 times with random initialization, UniPAD and CLAP.
+Mean and standard deviation of mAP are $48.55\pm 0.18\%$ (Rand.), $49.66\pm 0.29\%$ (UniPAD) and $51.16\pm 0.10\%$ (CLAP).
+CLAP achieves the best average performance and robustness against random seeds.
+B.3 Transferring to Other Datasets
+We conduct further experiments to evaluate the transferring ability of CLAP.
+Specifically, we select LiDAR-based 3D object detector CenterPoint Yin et al.
+(2021) on Once Mao et al.
+(2021) for the downstream task.
+A 40-beam LiDAR is utilized in Once Mao et al.
+(2021) to collect 15k labeled training data.
+We randomly sample $5\%$ and also use all of the labeled training set to train the from-scratch model until convergence is observed.
+Then we use pre-trained weights by CLAP on NuScenes Caesar et al.
+(2020) to initialize the same model and fine-tune it with the same training iterations as the randomly initialized model.
+Results are shown in Table 5.
+It can be found that pre-training by CLAP also benefits LiDAR-based 3D object detection, even in a cross-dataset setting.
+And if we look at the performance of “Rand*” and “CLAP*”, CLAP also accelerates the convergence in downstream task.
+Init. F.T. mAP Vehicle Pedestrian Cyclist
+0-30m 30-50m 50m- 0-30m 30-50m 50m- 0-30m 30-50m 50m-
+Rand* 5% 20.48 58.03 25.22 12.98 11.62 9.75 6.97 21.55 6.83 3.11
+CLAP* 22.86 +2.38 58.37 26.38 14.07 12.60 9.50 7.88 30.08 10.50 5.39
+Rand 46.07 76.71 51.15 31.84 37.53 20.12 9.84 62.00 42.61 24.18
+CLAP 46.88 +0.81 76.98 51.64 31.31 38.79 20.60 9.74 63.75 43.21 26.83
+Rand* 100% 64.00 86.21 70.20 58.20 57.80 41.18 23.55 75.95 61.45 45.80
+CLAP* 64.74 +0.74 88.14 72.59 59.13 57.37 42.24 24.22 77.11 61.91 45.63
+Rand 65.03 88.18 74.23 61.75 57.32 38.90 21.96 78.07 64.32 48.16
+CLAP 65.56 +0.53 87.97 72.77 62.11 58.33 40.11 21.29 78.63 64.70 47.27
+Table 5: Results for transferring experiments on Once Mao et al.
+(2021) dataset.
+CenterPoint Yin et al.
+(2021) is used as the downstream detector.
+“Init.” indicates the initialization methods.
+“F.T.” indicates the number of training samples in fine-tuning stage.
+Overall mAP and APs for different categories within different ranges are shown in this table.
+“Rand*” means training the randomly initialized model with the original training iterations in OpenPCDet Team (2020).
+“Rand” indicates that we increase the number of training iterations for randomly initialized model until convergence is observed.
+“CLAP*” indicates that we pre-train the backbones with CLAP on NuScenes Caesar et al.
+(2020) and then fine-tune on Once with the original iterations in Team (2020).
+“CLAP” uses the same fine-tuning iterations as “Rand”.
+We use green color to highlight the performance improvement brought by CLAP.
+All the results are in %.
+Init. mAP NDS mAP mAPH
+Rand. 64.57 67.61 69.00 67.57
+UniPAD 64.69 68.06 69.03 67.56
+CLAP 65.11 68.38 69.49 68.04
+Table 6: Fine-tuning on $100\%$ NuScenes and Waymo with fixed random seed.
+B.4 More Fine-tuning Data
+We fine-tune on $100\%$ NuScenes and Waymo with fixed random seed to provide more insights.
+Here, we show mAP and NDS for NuScenes in $2^{nd}$ and $3^{rd}$ cols and Level-2 mAP, mAPH for Waymo.
+CLAP improves random initialization by $\color[rgb]{0.1328125,0.546875,0.1328125}{\mathbf{+0.55}}$} whereas UniPAD by $\color[rgb]{0.1328125,0.546875,0.1328125}{\mathbf{+0.15}}$} on average, showing CLAP’s effectiveness.
+With Table 3 in the paper, more gain can be expected with more unlabeled data in real scenario.
+Appendix C The Use of Large Language Models.
+We use Large Language Models to help and aid editing/polishing the paper.
+In details, we polish the Introduction and Related Work sections with Large Language Models, which mostly focuses on grammar, spelling and word choice.